### v1

# ULTra: Unveiling Latent Token Interpretability in Transformer-Based Understanding

Abstract
Transformers have revolutionized Computer Vision (CV) and Natural Language Processing (NLP) through self-attention mechanisms. However, due to their complexity, their latent token representations are often difficult to interpret. We introduce a novel framework that interprets Transformer embeddings, uncovering meaningful semantic patterns within them. Based on this framework, we demonstrate that zero-shot unsupervised semantic segmentation can be performed effectively without any fine-tuning using a model pre-trained for tasks other than segmentation. Our method reveals the inherent capacity of Transformer models for understanding input semantics and achieves state-of-the-art performance in semantic segmentation, outperforming traditional segmentation models. Specifically, our approach achieves an accuracy of 67.2 % and an mIoU of 32.9 % on the COCO-Stuff dataset, as well as an mIoU of 51.9 % on the PASCAL VOC dataset. Additionally, we validate our interpretability framework on LLMs for text summarization, demonstrating its broad applicability and robustness.
1 Introduction
In recent years, the Transformer architecture and foundation models, leveraging self-attention mechanisms to capture complex dependencies in text, have transformed Natural Language Processing (NLP) benchmarks [58, 54, 2, 51]. Similarly, Vision Transformers (ViTs) [13] have been adapted in Computer Vision (CV) and now serve as the backbone for various tasks such as segmentation and object detection [52, 30]. Despite their success, understanding the interpretability of Transformers remains a challenge due to the complexity of their latent token representations.
${\zeta=0.5}$${\zeta=0.35}$${\zeta=0.2}$
Segmentation Results
Original Image
Figure 1: Hierarchical clustering tree showing the grouping of token relevance maps for all tokens in a latent layer of the Vision Transformer, not limited to the CLS token. Each leaf node represents a single token relevance map, while higher-level nodes show aggregated clusters based on a clustering threshold $(\zeta)$, which controls the level of detail. Lower $\zeta$ values reveal finer details, while higher values create broader, more general clusters. This approach demonstrates how pre-trained Vision Transformers can perform unsupervised semantic segmentation, identifying meaningful patterns within token representations without requiring additional training or fine-tuning.
Several methods have been developed to enhance the interpretability of CNN-based models [47, 63, 44]. While some of these can be extended to Transformer architectures, they do not fully leverage the unique attention mechanisms inherent to Transformers. Recent research has introduced interpretability methods specifically designed for Transformers [6, 1, 59]. However, these approaches primarily focus on explaining final model outputs, providing limited insight into the intermediate processes that lead to predictions. For instance, [7] maps latent tokens into CLIP’s [37] multi-modal space to find corresponding text descriptions, relying on an external text encoder for interpretability. In contrast, our approach directly interprets the latent space of ViTs, elucidating the role and function of each token within the high-dimensional space without relying on external models.
This paper introduces a framework to interpret latent tokens, offering a deeper understanding of the internal workings of Transformers. This understanding enables users to perform image semantic segmentation using pre-trained Transformer-based vision models in an unsupervised, zero-shot manner, without any additional training. We demonstrate that applying semantic segmentation based on our interpretability framework achieves state-of-the-art performance on benchmark image segmentation datasets.
Drawing inspiration from [6], our method analyzes the semantic information retained by latent tokens, enabling tasks such as object selection and semantic segmentation without additional training. We demonstrate that Transformers inherently understand the semantic structure of their input, viewing it as a collection of distinct concepts. Each latent token identifies a specific concept with semantic significance, thereby shedding light on the decision-making process of these models.
As shown in Section 4, our framework proves effective in a range of tasks, including semantic segmentation and model interpretation. Most recent unsupervised semantic segmentation methods involve an additional training phase to learn feature representations [46, 18, 27]. Our approach, however, utilizes the understanding embedded in pre-trained models to achieve zero-shot segmentation, leveraging their inherent knowledge of images. The stronger a model’s comprehension of image content, the more accurately it performs segmentation. We further demonstrate that our method is capable of interpreting large language models (LLMs) at the token level, validating its application in tasks such as text summarization. The main contributions of this paper are as follows:
  • We propose a framework for interpreting latent tokens in Transformers, providing valuable insights into model decision-making processes.
  • By aggregating relevance maps generated by tokens using hierarchical clustering, we achieve zero-shot unsupervised semantic segmentation on pre-trained models, outperforming SOTA methods that require additional training.
  • We demonstrate the capability of our method to interpret LLMs at the token level, showcasing its practical applicability to textual data. Specifically, we demonstrate the interpretation of LLM operations in the text summarization task.
This paper is structured as follows: Section 2 reviews interpretability frameworks and previous work on semantic segmentation, highlighting our SOTA results. Section 3 presents our interpretability framework. In Section 4, we showcase its effectiveness on image and text tasks. Finally, Section 5 concludes the paper.
2 Related Work
The interpretability of deep learning architectures has become a central focus in AI research [16]. As models grow in complexity, understanding their decision-making processes is essential for ensuring transparency, reliability, and fairness [53]. Interpretability not only aids in debugging and performance improvement but also builds trust in AI systems, particularly in fields like healthcare, finance, and autonomous driving. Opaque models can perpetuate biases and generate unforeseen outcomes, making interpretability crucial for bridging high performance with safe, practical AI deployment [26]. In Transformer models, tokens are key to interpreting behavior, with their relationship to spatial locations in images or sequence order adding an important layer to interpretation. This relationship enhances semantic segmentation as a measure for evaluating token-based interpretation frameworks, which this section will explore.
2.1 Model Interpretation
Model interpretability is a critical area of research in deep learning, especially for complex models like transformers. Traditional deep learning models, such as CNNs, have been interpreted using various techniques like saliency maps [47], deconvolutional networks [63], Guided Backpropagation [48], Feature Visualization [34], Local Interpretable Model-Agnostic Explanations (LIME) [40], SHapley Additive exPlanations (SHAP) [31], Class Activation Mapping (CAM) [65], Feature Attribution with Integrated Gradients [50], and Grad-CAM [44], which highlight the most important regions of the input that influence model predictions. While effective for CNNs, these techniques are less suitable for transformer architectures, as they fail to account for the unique self-attention mechanisms that transformers rely on. Anchors is a model-agnostic interpretability method that provides high-precision, locally faithful explanations by generating if-then rules that sufficiently explain a model’s prediction for specific inputs [41].
Recent interpretability research on transformers has focused on understanding how these models allocate attention and propagate information. A seminal contribution was made by examining self-attention patterns in transformer-based language models, revealing how attention is distributed across tokens [59]. Abnar et al. [1] advanced this by proposing methods to visualize attention flow across layers, aiding in the understanding of information propagation. Chefer et al. [6] introduced a relevance-based approach, computing relevance scores for each token to provide deeper insights into the model’s decision-making process. Additionally, the sensitivity of model predictions to input tokens has been explored as another way to interpret transformer behavior [20].
Interpretability methods for reflecting semantic relations within input sequences have been investigated in NLP tasks, aiming to characterize how transformers associate words or tokens across different positions in a sentence for downstream tasks [11]. Justifying the role of the multi-head attention mechanism in transformers and its contribution to improved performance is another area of study, revealing insights into how transformers operate [32]. Other notable interpretability methods include Interpretability-Aware Visual Transformers (IA-ViT) [36], Mechanistic Interpretability [39], Concept Transformers [42], Nested Hierarchical Transformers [64], Tracr [29], and ViT-net [22].
While many methods focus on explaining model outputs, fewer efforts have been directed toward understanding the intermediate processes within the model. The internal representations, particularly the latent tokens in Vision Transformers (ViTs), remain largely unexplored. A similar approach is the work by [7], which focuses on the CLIP model [37]. Their method involves disabling the self-attention mechanism to map the latent tokens into a multimodal space without additional training. However, simply disabling self-attention may introduce distribution shifts, raising concerns about the validity of their results. Furthermore, their approach relies heavily on CLIP’s text encoder, which limits its generalizability to other models. Our work overcomes these limitations by focusing on backward-looking analysis, investigating how latent tokens relate to the input rather than looking ahead. Additionally, we rely solely on one encoder, making our method more applicable to a broader range of models.
2.2 Unsupervised Semantic Segmentation
Unsupervised semantic segmentation has advanced significantly due to self-supervised learning and clustering-based techniques, which reduce reliance on labeled datasets. Early approaches, such as Invariant Information Clustering (IIC) [21], employed mutual information to group similar pixels without using labels. Building on this foundation, PiCIE introduced consistency by integrating photometric and geometric invariances, setting an important precedent for subsequent methods [9].
The emergence of Vision Transformers (ViTs) and self-supervised learning marked a major shift in the field. DINO became a pioneering method, extracting rich, meaningful features without the need for labeled data [5]. Its effectiveness in unsupervised segmentation laid the groundwork for more recent advancements. For instance, TransFGU [62] performs semantic segmentation in a top-down manner by deriving class activation maps from DINO models. STEGO [18] leverages DINO’s features, employing contrastive learning to group similar regions and achieving notable improvements in segmentation accuracy.
Subsequent methods further refined these concepts. MaskContrast incorporated clustering techniques to ensure region consistency across different views, enhancing feature representations for unsupervised segmentation [56]. Leopart utilized self-supervised ViTs to improve pixel grouping, particularly in complex scenes [66]. ACSeg introduced adaptive conceptualization for pixel-level semantic grouping, using an Adaptive Concept Generator (ACG) to dynamically align learnable prototypes with relevant concepts, thereby addressing over- and under-clustering challenges in varied images [27].
Leveraging hidden positives for unsupervised semantic segmentation enhances pixel grouping by identifying and utilizing implicit positive relationships within the data, boosting segmentation performance without the need for labeled examples [45]. Using self-supervised learning and adaptive feature representations to enable models to discover and segment novel, unseen object categories without labeled data have also been investigated [55]. Smooseg introduces a smoothness prior to enhancing unsupervised semantic segmentation by promoting coherent region grouping [25]. Unsupervised semantic segmentation is also improved by leveraging depth-guided feature correlation and targeted sampling to enhance region consistency and accuracy [46]. [17].
Unlike approaches that depend on self-training, pseudo-labeling, or complex setups, our model ULTra introduces a zero-shot method for unsupervised semantic segmentation. ULTra achieves strong segmentation performance by directly leveraging the semantic information embedded in the latent tokens of pre-trained Transformer models, without the need for additional training or fine-tuning. Furthermore, ULTra emphasizes explainability within Transformers by illustrating how the model’s latent tokens contribute to segmentation, in contrast to the common practice of using only the CLS token to represent the entire image. This zero-shot, explainability-driven approach offers a novel direction for unsupervised segmentation, demonstrating that Transformer-based architectures can achieve SOTA performance without labeled data or extensive additional processing.
(a) Original Image         (b) Latent Token’s Relevance Map (c) Predicted Binary Mask
Refer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 2: An example of token interpretation by our model and its predicted binary mask. (a) Original image. (b) Overlay of $\tilde{S}_{i}^{(13)}$ on the original image for different $i$, where the location of the $i$-th token is indicated by the purple square. (c) The binary mask $M_{i}^{(13)}$ for each corresponding relevance map in (b).
3 Methodology
In this section, we present our approach for interpreting latent representations in Transformers. We begin with essential preliminaries and notation for clarity, followed by a detailed explanation of our method for interpreting latent tokens and its application to semantic segmentation.
3.1 Preliminaries and Notation
The architecture of a typical Transformer can be formulated as follows: the input $X$ is split into $n$ tokens $\{\mathbf{x}_{i}\}_{i=1}^{n}$. After tokenization, token embeddings $\{\mathbf{e}_{i}\}_{i=0}^{n}$ are computed, where $\mathbf{e}_{0}$ corresponds to the CLS token. Positional encodings $\text{PE}_{i}$ are added to the $i$-th token embedding to incorporate spatial information, resulting in the latent token representation $\mathbf{z}_{i}^{(1)}=\mathbf{e}_{i}+\text{PE}_{i}$. Here, $\mathbf{z}_{i}^{(l)}$ represents a latent token, where $l$ denotes the layer index with $l\in\{1,\dots,L\}$ and $L$ is the total number of layers in the Transformer, and $i$ represents the $i$-th token within the $l$-th layer.
For each head $h\in\{1,\dots,H\}$ in the multi-head attention mechanism, the queries, keys, and values corresponding to the $i$-th token are obtained via linear transformations, projecting the latent token of dimension $d$ into dimension $k$:
$Q^{(l)}_{h}(\mathbf{z}_{i}^{(l-1)})=(W_{h,q}^{(l)})^{T}\mathbf{z}_{i}^{(l-1)},K^{(l)}_{h}(\mathbf{z}_{i}^{l-1})=(W_{h,k}^{(l)})^{T}\mathbf{z}^{(l-1)}_{i},$
$V^{(l)}_{h}(\mathbf{z}_{i}^{(l-1)})=(W_{h,v}^{(l)})^{T}\mathbf{z}_{i}^{(l-1)},\quad\forall l\in\{2,\dots,L\}$ (1)
where $W_{h,q}^{(l)},W_{h,k}^{(l)},W_{h,v}^{(l)}\in\mathbb{R}^{d\times k}$. The attention weights for each token pair $(i,j)$ at layer $l$ and head $h$ are computed as:
$\alpha_{h,i,j}^{(l)}=\text{softmax}_{j}\left(\frac{\langle Q^{(l)}_{h}(\mathbf{z}_{i}^{l-1}),K^{(l)}_{h}(\mathbf{z}_{j}^{(l-1)})\rangle}{\sqrt{k}}\right).$ (2)
Then, $i$-th token is updated by summing over the weighted values across all heads:
$\mathbf{\bar{u}}^{(l)}_{i}=\sum_{h=1}^{H}(W_{c,h}^{(l)})^{T}\sum_{j=1}^{n}\alpha_{h,i,j}^{(l)}V^{(l)}_{h}(\mathbf{z}_{j}^{(l-1)}),$ (3)
where $W_{c,h}\in\mathbb{R}^{k\times d}$. The updated token representation $\mathbf{u}_{i}$ after the attention layer is computed as:
$\mathbf{u}_{i}^{(l)}=\text{LayerNorm}(\mathbf{z}_{i}^{(l-1)}+\mathbf{\bar{u}}^{(l)}_{i}).$ (4)
Each token then passes through a feed-forward network:
$\mathbf{\bar{z}}_{i}^{(l)}$ $=(W_{2}^{(l)})^{T}\text{ReLU}((W_{1}^{(l)})^{T}\mathbf{u}_{i}),$ (5)
$\mathbf{z}_{i}^{(l)}$ $=\text{LayerNorm}(\mathbf{u}_{i}+\mathbf{\bar{z}}_{i}^{(l)}).$ (6)
Here, $W_{1}^{(l)}\in\mathbb{R}^{d\times m}$, $W_{2}^{(l)}\in\mathbb{R}^{m\times d}$.
3.2 Interpreting Latent Tokens
A straightforward approach for interpreting a model involves analyzing the semantic flow from the input to the corresponding logits. This can be achieved by adding the attention probability matrix to an identity matrix $I$, which incorporates skip connections, and then multiplying the result across layers [1]. However, a notable challenge arises from multiple attention heads in each Transformer layer. To address this, [6] proposes performing a weighted average across attention heads, with the weights determined by the gradient of the logits with respect to the attention weights. We employed a modified, simpler version of this method, utilizing only the attention weights rather than the layer’s relevance.
In our framework, latent tokens $\mathbf{z}_{i}^{(l)}$ are represented as vectors rather than direct class correspondences, introducing an additional layer of complexity. To manage this, an appropriate transformation, such as the energy or euclidean norm of the latent token, is employed as a surrogate. Consequently, we compute the gradient of $\|\mathbf{z}_{i}^{(l)}\|$ with respect to the attention weights to facilitate the analysis.
In the following equations, we define the relevance map $S_{i}^{(l)}\in\mathbb{R}^{n}$, where the $j$-th element of this vector represents the importance of the $j$-th input token to the targeted latent token $\mathbf{z}_{i}^{(l)}$. The relevance map is computed as:
$\overline{\mathbf{A}}_{i}^{(b,l)}$ $=I+\mathbb{E}_{h}\left(\nabla_{\mathbf{A}^{b}_{h}}\lVert\mathbf{z}_{i}^{(l)}\rVert\odot\mathbf{A}^{(b)}_{h}\right)^{+},$
$\overline{S}_{i}^{(l)}$ $=\overline{\mathbf{A}}_{i}^{(1,l)}\cdot\overline{\mathbf{A}}_{i}^{(2,l)}\cdot\;\dots\;\cdot\overline{\mathbf{A}}_{i}^{(l-1,l)},$
$S_{i}^{(l)}$ $=\overline{S}_{i}^{(l)}[i,1:\;],$ (7)
where $(\cdot)^{+}$ means considering only positive values, $S,\;\overline{\mathbf{A}}_{i}^{(b,l)}\in\mathbb{R}^{(n+1)\times(n+1)}$, $\odot$ denotes the Hadamard product, $\mathbb{E}_{h}$ represents the mean across the heads dimension, and $\mathbf{A}^{(b)}_{h}\in\mathbb{R}^{h\times(n+1)\times(n+1)}$ is the attention score matrix in the $b$-th layer and $A_{h,i,j}^{(b)}=\alpha_{h,i,j}^{(b)}$.
Due to the skip connections in the transformer, most of the contribution of $S_{i}^{(l)}$ is concentrated on $S_{i}^{(l)}[i-1]$ which makes it hard to analyze other tokens’ contribution. To address this issue, we replace this element with the maximum value of other elements, thereby capturing the contributions of additional tokens to the selected token.
For vision tasks, we first reshape the relevance map and then upsample it using bilinear or cubic interpolation to match the resolution of the model’s input. The resulting higher-dimensional matrix is denoted as $\tilde{S}$. This upsampling step is essential for enabling accurate object selection and semantic segmentation tasks.
3.3 ULTra in Unsupervised Tasks
In this section, we aim to examine ULTra’s capability to adapt to various tasks involving semantic knowledge. Importantly, it requires no additional training, leveraging the inherent understanding within transformers rather than relying on loss functions objective, final layer outputs, or fine-tuning.
Unsupervised Semantic Segmentation. As previously discussed, a relevance map can be defined for each latent token at a fixed layer, with the total number of relevance maps equal to the number of latent tokens. In the context of segmentation, the goal is to assign a class label to each pixel within an image. To achieve this, we employ clustering techniques, such as hierarchical clustering, that do not require a predefined number of classes. These techniques group the relevance maps into $k$ distinct clusters, where $k$ is unknown. Ideally, we aim for $k$ to approximate the actual number of classes present in the image.
Our approach provides flexibility in adjusting the value of $k$ by modifying the cutoff distance threshold $\zeta$ within the clustering algorithm. Increasing $\zeta$ produces fewer, broader clusters that capture general categories, such as background and foreground. Conversely, reducing $\zeta$ allows for finer segmentation, distinguishing more specific features, such as an object’s head or hands. To prevent the method from disproportionately favoring larger objects, given that the number of elements in each cluster may vary, we apply min-max scaling to each cluster independently.
After clustering, we define $k$ distinct concepts by aggregating the relevance maps within each cluster. The aggregated relevance map $\tilde{S}^{(l)}_{c}[x,y]$ for a cluster $c$ is computed as:
$\tilde{S}^{(l)}_{c,\zeta}[x,y]=\sum_{i\in\phi_{\zeta}(c)}\tilde{S}_{i}^{(l)}[x,y],$ (8)
where $\phi_{\zeta}(c)=\{i\mid\text{Class}(\tilde{S}_{i}^{(l)})=c\}$ represents the set of label assignments determined by the clustering algorithm. For each pixel at position $[x,y]$, the class label is determined by identifying the cluster $c$ with the highest relevance value at that pixel. Mathematically, the class assignment for a pixel is expressed as:
$\text{Class}[x,y]_{\zeta}^{(l)}=\underset{c\in\{1,\ldots,k\}}{\text{argmax}}\,S_{c,\zeta}^{(l)}[x,y].$ (9)
Some examples illustrating our segmentation method is presented in Figure 3. Additionally, the hierarchy tree and the effect of the threshold are demonstrated in Figure 1.
Figure 3: ULTra segmentation results on sample images. The top row displays the original images, the middle row shows ground-truth annotations, and the bottom row presents our model’s predictions.
4 Experiment
Datasets. In our experiments, we evaluate model performance on several semantic segmentation benchmarks, focusing on vision-related tasks. We conducted experiments on three datasets: COCO-Stuff 27 [4], PASCAL VOC 2012 [15], and Potsdam-3 [19]. This combination of datasets provides a diverse testing ground to rigorously evaluate our unsupervised zero-shot approach across both standard and challenging perspectives in semantic segmentation.
COCO-Stuff 27, a subset of the COCO dataset [28], includes complex, real-world scenes with pixel-level annotations across various object categories. Similarly, PASCAL VOC 2012 serves as a classic benchmark with pixel-level annotations, while the Potsdam-3 dataset offers a unique aerial, top-down perspective of urban scenes. The latter adds an additional challenge with its high-resolution images of buildings, roads, and natural elements captured over the city of Potsdam.
For our qualitative analysis of LLM interpretation in the task of text summarization, as described in Section 4.3, we utilized the TL;DR dataset [49]. The TL;DR dataset contains summary comparisons with human feedback collected by OpenAI. Each entry consists of a Reddit post, including its title, original content, and a human-generated TL;DR.
Models. For all experiments in the vision tasks, we used CLIP’s image encoder ViT-B/32 [37]. For interpreting text summarization, as described in Section 4.3, we used the Llama-2-7B language model [54]. All experiments were run on 8 NVIDIA RTX 4090-24GB GPUs.
4.1 Zero-shot Unsupervised Object Selection
The upsampled relevance map $\tilde{S}_{i}^{(l)}$ can be converted into a binary segmentation mask using a threshold $\tau$, where the binary mask $M_{i}^{(l)}$ is defined as:
$M_{i}^{(l)}[x,y]=\begin{cases}0,&\text{if }\tilde{S}_{i}^{(l)}[x,y]<\tau,\\ 1,&\text{otherwise}.\end{cases}$ (10)
Here, $\tilde{S}_{i}^{(l)}[x,y]$ represents the relevance value at position $[x,y]$ in $\tilde{S}_{i}^{(l)}$, with $\tau$ as the threshold. When $\tilde{S}_{i}^{(l)}[x,y]$ is below $\tau$, $M_{i}^{(l)}[x,y]$ is set to 0; otherwise, it is set to 1, marking the object region.
Our findings indicate that as tokens propagate through the network, they refine their object representation while retaining the semantic meaning of their associated image patches performing Object Selection. Figure 5 visually illustrates this process. For a given patch token $\mathbf{x}_{i}$, the object it most strongly represents is denoted as class $k_{i}$, indicating that $\mathbf{x}_{i}$ predominantly corresponds to a region of the object belonging to class $k_{i}$ in the image. The latent token $\mathbf{z}_{i}^{(l)}$ generates a relevance map that highlights areas with higher values associated with class $k_{i}$. After applying a threshold, this map becomes a binary segmentation mask expected to exhibit a high Intersection over Union (IoU) with the corresponding class $k_{i}$ region in the image. An illustrative example is shown in Figure 2.
To quantify alignment, we compute the IoU by converting the relevance map $S_{i}^{(l)}$ into a binary mask $M_{i}^{(l)}$ and comparing it with the ground-truth mask. We propose the Initial Token IoU (ITIoU) metric, which measures how well the relevance maps of input tokens align with their respective class masks. The ITIoU is calculated as:
$\textit{ITIoU}^{(l)}(X)=\frac{1}{C}\sum_{i=1}^{C}\frac{1}{|\mathcal{P}_{i}|}\sum_{\mathbf{x}_{j}\in\mathcal{P}_{i}}\text{IoU}(M_{j}^{(l)},G_{i}),$ (11)
where $C$ denotes the number of classes, $\mathcal{P}_{i}$ represents the set of tokens associated with class $i$, $M_{j}^{(l)}$ is the binary segmentation mask for token $\mathbf{x}_{j}$ within class $i$, and $G_{i}$ is the ground-truth mask for class $i$ in image $X$. The inner sum averages the IoU for tokens in $\mathcal{P}_{i}$ for each class, and the outer sum then averages across all classes. Using a threshold of 0.2, our ITIoU metric achieves an average score of 37.84 % on the COCO validation dataset and 39.51 % on the VOC dataset.
The ITIoU metric provides a comprehensive evaluation by incorporating a weighted average across tokens in each class, enhancing the assessment of token alignment with their respective ground-truth labels.
Refer to caption Refer to caption
(a) TL;DR: I’ve had very bad luck with girls my whole (b) TL;DR: I still like her but my rational
life and I don’t know how to get my confidence up. side says ”no, she is a trash person”.
Figure 4: Visualization of Token Contribution Scores ($\lambda_{i}^{(l)}$) highlighting the relevance of context tokens in interpreting the summary. Each token is colored proportionally to its $\lambda_{i}^{(l)}$ value. These visualizations demonstrate the model’s ability to identify key semantic elements in the context for generating relevant summaries.
4.2 Zero-shot Unsupervised Semantic Segmentation
We benchmarked the segmentation capability of our method against several approaches in the literature on unsupervised segmentation. Notably, unlike other methods, our approach requires no additional training. Instead, it relies solely on a pre-trained vision transformer, which may have been trained on tasks unrelated to unsupervised segmentation.
To evaluate our method, we used the Unsupervised mean Intersection over Union (U. mIoU) and Unsupervised pixel Accuracy (U. Accuracy) metrics, following standard practices in semantic segmentation research. In all experiments across datasets, as shown in Tables 1, 2, and 3, we set the cutoff distance threshold $\zeta=0.4$.
Table 1: Comparison of unsupervised segmentation methods on the 27-class COCO-Stuff validation set. ”W” indicates methods requiring additional training, while ”W/O” denotes methods without additional training.
Method Training U. Accuracy U. mIoU
IIC [21] W 21.8 6.7
DINO [5] W 30.5 9.6
PiCIE [9] W 48.1 13.8
ACSeg [27] W - 16.4
STEGO [18] W 56.9 28.2
DepthG [46] W 58.6 29.0
U2Seg [33] W 63.9 30.2
ULTra (Ours) W/O 67.2 32.9
Table 2: Comparison of unsupervised segmentation methods on the PASCAL VOC validation set.
Method Training U. mIoU
IIC [21] W 9.8
MaskContrast [56] W 35.0
Leopart [66] W 41.7
TransFGU [62] W 37.2
MaskDistill [57] W 42.0
ACSeg [27] W 47.1
ULTra (Ours) W/O 51.9
Table 3: Comparison of unsupervised segmentation methods on the Potsdam validation set.
Method Training U. ACC
IIC [21] W 65.1
DINO [5] W 71.3
STEGO [18] W 77.0
DepthG [46] W 80.4
ULTra (Ours) W/O 74.6
Our results in Tables 1, 2, and 3 highlight the effectiveness of the proposed method for unsupervised semantic segmentation. Unlike baseline approaches, which require additional training specifically tailored to segmentation, our method leverages a pre-trained vision transformer without any further fine-tuning. Despite this lack of task-specific training, our approach consistently outperforms other methods, demonstrating adaptability across diverse datasets.
4.3 Interpreting LLMs in Text Summarization
In this section, we examine how our interpretability framework can be applied to text summarization tasks to uncover the underlying mechanisms and intent of LLMs. By visualizing the regions of the input context that an LLM prioritizes while interpreting a given TL;DR summary, we gain deeper insights into the model’s behavior and decision-making processes. As shown in Figure 4, these visualizations allow us to pinpoint the most influential regions of a textual input prompt in generating concise and relevant summaries.
For the experiments, we used a Supervised Fine-Tuned (SFT) version of Llama-2-7B trained on the UltraFeedback Binarized (UFB) dataset [12]. Additionally, we aligned the model to the text summarization task on the TL;DR dataset using the Direct Preference Optimization (DPO) method [38] for 1,000 iterations, with a learning rate of $5\times 10^{-6}$ and $\beta=0.5$. To validate our framework, we selected the preferred response (TL;DR) of each sample in the dataset, denoted it as $y$, and used it as the summary of the context $x$.
In this task, we concatenate the context $x$ and the summary $y$ with a separator token. After feeding this input to the model, we compute the relevance scores of the TL;DR tokens with respect to the context tokens. We then average these scores for each token in $x$ to obtain a scalar value, referred to as the Token Contribution Score, $\lambda_{i}^{(l)}\in\mathbb{R}^{+}$, which highlights the contribution of each context token in interpreting the summary $y$ with respect to the context. This provides visual evidence of how the model processes the context text and identifies key semantic elements relevant to producing the summary $y$. Accordingly, $\lambda_{i}^{(l)}$ is computed as:
$\lambda_{i}^{(l)}=\frac{1}{|y|}\sum_{j=1}^{|y|}S_{j+|x|}^{(l)}[i],\quad\forall i\in\{1,\cdots,|x|\},$ (12)
where ||$\cdot$|| ⋅ | denotes the number of tokens in the text.
(a) Original Image (b) Explanation Relevency Map
Refer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 5: An example illustrating the model’s decision-making process across layers. Moving from left to right corresponds to deeper layers in the network. The first row corresponds to the CLS token, while the second, third, and fourth rows represent three different tokens, highlighted by red squares.
(a) COCO-Stuff
(b) PASCAL VOC
Figure 6: Ablation study on two evaluation metrics across layers. These plots demonstrate a progressive improvement in semantic segmentation performance in the deeper layers of the transformer model. This enhancement is attributed to latent tokens capturing more meaningful segment structures, resulting in increasingly accurate and refined semantic representations.
In example (a): semantically significant words such as ‘relationship’, ‘experience’, ‘rejection’, and ‘never’ are prominently highlighted, reflecting the model’s interpretation of the person’s struggles with relationships and feelings of rejection. Additionally, the highlighting of the question at the beginning of the context ‘How do I stop feeling bad…’ suggests the model recognizes the presence of uncertainty and a request for guidance, which is encapsulated in the summary as ‘I don’t know.’
In example (b): $\lambda_{i}^{(l)}$ scores reveals the model’s focus on words such as ‘feelings’, ‘hate’, ‘disappoint’, ‘love’, and ‘like’, which correspond to the person’s mixed emotions toward their girlfriend, as described in the summary. The apparent contradiction between ‘love’ and ‘trashness’ in the summary seems to be derived from these highlighted terms, suggesting the model understands the conflicting emotions present in the text. Furthermore, the emphasis on ‘character’ aligns with the summary’s judgmental tone, suggesting that the model interprets this word as indicative of an assessment of personality traits and behaviors.
This token-level analysis can serve as a valuable tool for effective supervision in future research, particularly for developing interpretable fine-tuning and alignment methods like Reinforcement Learning from Human Feedback (RLHF) [10, 49, 35] and Preference Optimization [38, 3, 14], where understanding model behavior and intent is essential. It may also provide useful insights in related areas, such as Chain-of-Thought [8, 24, 43] and Theory-of-Mind [60, 23, 61], which seek to make the reasoning processes and intentions of LLMs more transparent. Ultimately, our framework offers insights into fundamental questions in NLP, such as: ”How is the model thinking?” or ”What is the underlying intent behind the model’s generated response?”.
4.4 Ablation Study on the Effect of Layer Depth in ViT Token Understanding
In this section, we analyze the impact of depth on our model’s interpretability and segmentation performance, providing insights into the contribution of each layer. As anticipated and observed in Figure 6, deeper layers generally carry more semantic significance. However, the contribution diminishes in the final layers, suggesting that a depth of around 13 layers might be more than sufficient for the ViT to effectively comprehend image content. This finding implies that even fewer layers might achieve comparable results, potentially reducing computational costs without compromising performance.
We observe an intriguing behavior in the initial layers, where performance initially declines before improving. This phenomenon is also visually evident in Figure 5, where the attention maps in the first layer appear to focus on the entire image. This suggests that, initially, the token examines the image as a whole before selectively gathering information from tokens with similar characteristics.
5 Conclusion
This paper introduces a novel interpretability framework that provides valuable insights into the decision-making processes of Transformer models. The framework is based on the input interpretation provided by each token in an arbitrary layer. By aggregating the interpretation of all tokens in the same layer, the framework enhances understanding of the Transformer’s behavior at the layer level. The richness of this understanding enables zero-shot unsupervised semantic segmentation, where our method achieves state-of-the-art performance with an accuracy of $67.2\%$ and an mIoU of $32.9\%$ on the COCO-Stuff dataset, and an mIoU of $51.9\%$ on the PASCAL VOC dataset for the unsupervised zero-shot version. Beyond vision tasks, the framework is not limited to vision models and can also be applied to interpret the behavior of large language models (LLMs) in tasks such as text summarization. This highlights the versatility and broad applicability of the proposed framework across various domains, offering potential for future work to simultaneously demystify the multimodal understanding of Transformers.

### v2

@@
-ULTra: Unveiling Latent Token Interpretability in Transformer-Based Understanding
+ULTra: Unveiling Latent Token Interpretability in Transformer-Based Understanding and Segmentation
 Abstract
-Transformers have revolutionized Computer Vision (CV) and Natural Language Processing (NLP) through self-attention mechanisms.
-However, due to their complexity, their latent token representations are often difficult to interpret.
-We introduce a novel framework that interprets Transformer embeddings, uncovering meaningful semantic patterns within them.
-Based on this framework, we demonstrate that zero-shot unsupervised semantic segmentation can be performed effectively without any fine-tuning using a model pre-trained for tasks other than segmentation.
-Our method reveals the inherent capacity of Transformer models for understanding input semantics and achieves state-of-the-art performance in semantic segmentation, outperforming traditional segmentation models.
-Specifically, our approach achieves an accuracy of 67.2 % and an mIoU of 32.9 % on the COCO-Stuff dataset, as well as an mIoU of 51.9 % on the PASCAL VOC dataset.
-Additionally, we validate our interpretability framework on LLMs for text summarization, demonstrating its broad applicability and robustness.
+Transformers have revolutionized Computer Vision (CV) through self-attention mechanisms.
+However, their complexity makes latent token representations difficult to interpret.
+We introduce ULTra, a framework for interpreting Transformer embeddings and uncovering meaningful semantic patterns within them.
+ULTra enables unsupervised semantic segmentation using pre-trained models without requiring fine-tuning.
+Additionally, we propose a self-supervised training approach that refines segmentation performance by learning an external transformation matrix without modifying the underlying model.
+Our method achieves state-of-the-art performance in unsupervised semantic segmentation, outperforming existing segmentation methods.
+Furthermore, we validate ULTra for model interpretation on both synthetic and real-world scenarios, including Object Selection and interpretable text summarization using LLMs, demonstrating its broad applicability in explaining the semantic structure of latent token representations.
 1 Introduction
-In recent years, the Transformer architecture and foundation models, leveraging self-attention mechanisms to capture complex dependencies in text, have transformed Natural Language Processing (NLP) benchmarks [58, 54, 2, 51].
+In recent years, the Transformer architecture and foundation models, leveraging self-attention mechanisms to capture complex dependencies in text, have transformed Natural Language Processing (NLP) benchmarks [50, 47, 45].
 Similarly, Vision Transformers (ViTs) [17] have been adapted in Computer Vision (CV) and now serve as the backbone for various tasks such as segmentation and object detection [46, 30].
 Despite their success, understanding the interpretability of Transformers remains a challenge due to the complexity of their latent token representations.
-${\zeta=0.5}$${\zeta=0.35}$${\zeta=0.2}$
-Segmentation Results
-Original Image
-Figure 1: Hierarchical clustering tree showing the grouping of token relevance maps for all tokens in a latent layer of the Vision Transformer, not limited to the CLS token.
-Each leaf node represents a single token relevance map, while higher-level nodes show aggregated clusters based on a clustering threshold $(\zeta)$, which controls the level of detail.
-Lower $\zeta$ values reveal finer details, while higher values create broader, more general clusters.
-This approach demonstrates how pre-trained Vision Transformers can perform unsupervised semantic segmentation, identifying meaningful patterns within token representations without requiring additional training or fine-tuning.
 Several methods have been developed to enhance the interpretability of CNN-based models [43, 56, 40].
@@
 In contrast, our approach directly interprets the latent space of ViTs, elucidating the role and function of each token within the high-dimensional space without relying on external models.
-This paper introduces a framework to interpret latent tokens, offering a deeper understanding of the internal workings of Transformers.
-This understanding enables users to perform image semantic segmentation using pre-trained Transformer-based vision models in an unsupervised, zero-shot manner, without any additional training.
-We demonstrate that applying semantic segmentation based on our interpretability framework achieves state-of-the-art performance on benchmark image segmentation datasets.
-Drawing inspiration from [6], our method analyzes the semantic information retained by latent tokens, enabling tasks such as object selection and semantic segmentation without additional training.
-We demonstrate that Transformers inherently understand the semantic structure of their input, viewing it as a collection of distinct concepts.
-Each latent token identifies a specific concept with semantic significance, thereby shedding light on the decision-making process of these models.
-As shown in Section 4, our framework proves effective in a range of tasks, including semantic segmentation and model interpretation.
-Most recent unsupervised semantic segmentation methods involve an additional training phase to learn feature representations [46, 18, 27].
-Our approach, however, utilizes the understanding embedded in pre-trained models to achieve zero-shot segmentation, leveraging their inherent knowledge of images.
-The stronger a model’s comprehension of image content, the more accurately it performs segmentation.
-We further demonstrate that our method is capable of interpreting large language models (LLMs) at the token level, validating its application in tasks such as text summarization.
-The main contributions of this paper are as follows:
-  • We propose a framework for interpreting latent tokens in Transformers, providing valuable insights into model decision-making processes.
-  • By aggregating relevance maps generated by tokens using hierarchical clustering, we achieve zero-shot unsupervised semantic segmentation on pre-trained models, outperforming SOTA methods that require additional training.
-  • We demonstrate the capability of our method to interpret LLMs at the token level, showcasing its practical applicability to textual data.
-Specifically, we demonstrate the interpretation of LLM operations in the text summarization task.
-This paper is structured as follows: Section 2 reviews interpretability frameworks and previous work on semantic segmentation, highlighting our SOTA results.
-Section 3 presents our interpretability framework.
-In Section 4, we showcase its effectiveness on image and text tasks.
-Finally, Section 5 concludes the paper.
+In this paper, we introduce Unveiling Latent Token Interpretability in Transformer-Based Understanding (ULTra) a framework for interpreting latent tokens in Transformers.
+ULTra provides deeper insights into the internal representations of these models.
+By analyzing the semantic information captured by latent tokens, we show that Transformers inherently encode the semantic structure of their input as a collection of distinct concepts.
+Moreover, by aggregating explanation maps, ULTra achieves unsupervised semantic segmentation.
+Unlike prior unsupervised segmentation methods that require additional training [42, 20, 27], ULTra leverages the intrinsic knowledge of pre-trained models on tasks other than semantic segmentation to achieve state-of-the-art performance on benchmark datasets—without the need for fine-tuning.
+To further enhance segmentation performance, we introduce a self-consistency approach that learns an external transformation matrix in a self-supervised manner, refining segmentation without modifying the underlying model.
+Additionally, we validate our interpretability framework on transformer-based LLMs through qualitative analyses in text summarization, demonstrating its broad applicability across modalities.
+Our main contributions are as follows:
+  • We propose a framework for interpreting latent tokens in Transformers, uncovering the stored semantic knowledge.
+To the best of our knowledge, we are the first to directly investigate latent token interpretability.
+  • We extend ULTra to several designed synthetic and real-world tasks, showcasing its versatility and applicability, including Object Selection and interpretation of LLMs in text summarization tasks.
+  • By aggregating explanation maps generated by tokens, we achieve SOTA on pre-trained models, surpassing methods that require additional training.
+Furthermore, we further improve our segmentation by learning an external weight matrix through the proposed self-consistency loss.
 2 Related Work
-The interpretability of deep learning architectures has become a central focus in AI research [16].
-As models grow in complexity, understanding their decision-making processes is essential for ensuring transparency, reliability, and fairness [53].
-Interpretability not only aids in debugging and performance improvement but also builds trust in AI systems, particularly in fields like healthcare, finance, and autonomous driving.
-Opaque models can perpetuate biases and generate unforeseen outcomes, making interpretability crucial for bridging high performance with safe, practical AI deployment [26].
-In Transformer models, tokens are key to interpreting behavior, with their relationship to spatial locations in images or sequence order adding an important layer to interpretation.
-This relationship enhances semantic segmentation as a measure for evaluating token-based interpretation frameworks, which this section will explore.
-2.1 Model Interpretation
-Model interpretability is a critical area of research in deep learning, especially for complex models like transformers.
-Traditional deep learning models, such as CNNs, have been interpreted using various techniques like saliency maps [47], deconvolutional networks [63], Guided Backpropagation [48], Feature Visualization [34], Local Interpretable Model-Agnostic Explanations (LIME) [40], SHapley Additive exPlanations (SHAP) [31], Class Activation Mapping (CAM) [65], Feature Attribution with Integrated Gradients [50], and Grad-CAM [44], which highlight the most important regions of the input that influence model predictions.
-While effective for CNNs, these techniques are less suitable for transformer architectures, as they fail to account for the unique self-attention mechanisms that transformers rely on.
-Anchors is a model-agnostic interpretability method that provides high-precision, locally faithful explanations by generating if-then rules that sufficiently explain a model’s prediction for specific inputs [41].
-Recent interpretability research on transformers has focused on understanding how these models allocate attention and propagate information.
-A seminal contribution was made by examining self-attention patterns in transformer-based language models, revealing how attention is distributed across tokens [59].
-Abnar et al.
-[1] advanced this by proposing methods to visualize attention flow across layers, aiding in the understanding of information propagation.
-Chefer et al.
-[6] introduced a relevance-based approach, computing relevance scores for each token to provide deeper insights into the model’s decision-making process.
-Additionally, the sensitivity of model predictions to input tokens has been explored as another way to interpret transformer behavior [20].
-Interpretability methods for reflecting semantic relations within input sequences have been investigated in NLP tasks, aiming to characterize how transformers associate words or tokens across different positions in a sentence for downstream tasks [11].
-Justifying the role of the multi-head attention mechanism in transformers and its contribution to improved performance is another area of study, revealing insights into how transformers operate [32].
-Other notable interpretability methods include Interpretability-Aware Visual Transformers (IA-ViT) [36], Mechanistic Interpretability [39], Concept Transformers [42], Nested Hierarchical Transformers [64], Tracr [29], and ViT-net [22].
-While many methods focus on explaining model outputs, fewer efforts have been directed toward understanding the intermediate processes within the model.
-The internal representations, particularly the latent tokens in Vision Transformers (ViTs), remain largely unexplored.
-A similar approach is the work by [7], which focuses on the CLIP model [37].
-Their method involves disabling the self-attention mechanism to map the latent tokens into a multimodal space without additional training.
-However, simply disabling self-attention may introduce distribution shifts, raising concerns about the validity of their results.
-Furthermore, their approach relies heavily on CLIP’s text encoder, which limits its generalizability to other models.
-Our work overcomes these limitations by focusing on backward-looking analysis, investigating how latent tokens relate to the input rather than looking ahead.
-Additionally, we rely solely on one encoder, making our method more applicable to a broader range of models.
-2.2 Unsupervised Semantic Segmentation
-Unsupervised semantic segmentation has advanced significantly due to self-supervised learning and clustering-based techniques, which reduce reliance on labeled datasets.
-Early approaches, such as Invariant Information Clustering (IIC) [21], employed mutual information to group similar pixels without using labels.
-Building on this foundation, PiCIE introduced consistency by integrating photometric and geometric invariances, setting an important precedent for subsequent methods [9].
-The emergence of Vision Transformers (ViTs) and self-supervised learning marked a major shift in the field.
-DINO became a pioneering method, extracting rich, meaningful features without the need for labeled data [5].
-Its effectiveness in unsupervised segmentation laid the groundwork for more recent advancements.
-For instance, TransFGU [62] performs semantic segmentation in a top-down manner by deriving class activation maps from DINO models.
-STEGO [18] leverages DINO’s features, employing contrastive learning to group similar regions and achieving notable improvements in segmentation accuracy.
-Subsequent methods further refined these concepts.
-MaskContrast incorporated clustering techniques to ensure region consistency across different views, enhancing feature representations for unsupervised segmentation [56].
-Leopart utilized self-supervised ViTs to improve pixel grouping, particularly in complex scenes [66].
-ACSeg introduced adaptive conceptualization for pixel-level semantic grouping, using an Adaptive Concept Generator (ACG) to dynamically align learnable prototypes with relevant concepts, thereby addressing over- and under-clustering challenges in varied images [27].
-Leveraging hidden positives for unsupervised semantic segmentation enhances pixel grouping by identifying and utilizing implicit positive relationships within the data, boosting segmentation performance without the need for labeled examples [45].
-Using self-supervised learning and adaptive feature representations to enable models to discover and segment novel, unseen object categories without labeled data have also been investigated [55].
-Smooseg introduces a smoothness prior to enhancing unsupervised semantic segmentation by promoting coherent region grouping [25].
-Unsupervised semantic segmentation is also improved by leveraging depth-guided feature correlation and targeted sampling to enhance region consistency and accuracy [46].
-[17].
-Unlike approaches that depend on self-training, pseudo-labeling, or complex setups, our model ULTra introduces a zero-shot method for unsupervised semantic segmentation.
-ULTra achieves strong segmentation performance by directly leveraging the semantic information embedded in the latent tokens of pre-trained Transformer models, without the need for additional training or fine-tuning.
-Furthermore, ULTra emphasizes explainability within Transformers by illustrating how the model’s latent tokens contribute to segmentation, in contrast to the common practice of using only the CLS token to represent the entire image.
-This zero-shot, explainability-driven approach offers a novel direction for unsupervised segmentation, demonstrating that Transformer-based architectures can achieve SOTA performance without labeled data or extensive additional processing.
-(a) Original Image         (b) Latent Token’s Relevance Map (c) Predicted Binary Mask
+Interpretable Deep Learning.
+Model interpretability is a critical aspect of deep learning, particularly for complex architectures like Transformers.
+Traditional methods such as saliency maps [43], Grad-CAM [40], LIME [39], and SHAP [31] have been effective for CNNs but do not fully exploit Transformers’ self-attention mechanisms.
+A growing body of research focuses on Transformer-specific interpretability techniques that leverage attention mechanisms as intrinsic explanations [51, 1, 8, 22, 52].
+These methods are largely post-hoc, providing explanations after model training.
+In contrast, ante-hoc approaches such as IA-ViT [35] and Mechanistic Interpretability [38] seek to make Transformers inherently interpretable.
+Additionally, techniques like LeGrad [4] and IA-RED² [34] improve interpretability by analyzing feature formation and reducing redundancy in self-attention.
+Despite these advancements for model interpretability, understanding latent token in Vision Transformers (ViTs) remains underexplored, highlighting the need for methods that explicitly interpret latent tokens.
+A related study [9] explores latent token interpretation in CLIP by modifying self-attention mechanisms.
+Semantic Segmentation.
+Unsupervised semantic segmentation has progressed through self-supervised learning and clustering techniques.
+Early methods such as IIC [23] and PiCIE [11] leveraged mutual information and consistency principles to enhance feature representations.
+Transformer-based approaches like DINO [6] and STEGO [20] further improved segmentation by capturing meaningful structures through self-attention.
+Other techniques, including MaskContrast [48], Leopart [57], and ACSeg [27], refined segmentation through clustering and adaptive conceptualization.
+More recent methods, such as DepthG [42], incorporate depth-guided correlations, while U2SEG [32] utilizes pseudo-labeling.
+Additionally, SmooSeg [25] enforces smoothness priors, and HSG [24] applies hierarchical segmentation via multiview clustering Transformers.
+Moreover, in the context of semantic segmentation, Weakly Supervised Semantic Segmentation (WSSS) aims to generate segmentation masks using only image-level labels.
+These methods often rely on saliency maps from interpretability techniques, such as Class Activation Maps (CAMs), to localize objects [12, 54, 10].
+Typically, they depend on class logits to guide segmentation,refining masks based on predicted class scores.
+Latent Embedding.
+The high dimensionality and complex distribution of latent embeddings in deep models pose significant challenges for interpretation and manipulation.
+Methods like GroupViT[53] utilize hierarchical grouping to facilitate more meaningful representation learning, enabling semantic segmentation.
+Other approaches, such as [26, 3, 28], improve computational efficiency by eliminating redundant tokens, while register-based ViTs[16] address artifacts in feature maps caused by outlier-norm tokens through the introduction of register tokens.
+However, these techniques primarily emphasize computational efficiency and representation structuring rather than the interpretability of latent embeddings.
+[…]
+Figure 1: The overall architecture of the ULTra framework.
+The framework consists of a forward path, where the input data $\mathbf{x}$ is fed into the model, and a backward path, starting from the target layer $l$, where we compute the gradient of a scalar function of the $i$-th latent token, $f({\mathbf{z}_{i}^{(l)}})$, with respect to the attention probability matrix of the middle layer $b$.
+Next, we compute the corresponding contribution map $\mathbf{C}_{i}^{(b,l)}$ for all middle layers.
+Finally, we construct the explanation map $\overline{S}_{i}^{(l)}$, select its $i$-th row, and transform it to the input size.
+3 Methodology
+In this section, we introduce our approach for interpreting latent representations in Transformers.
+We start by outlining the essential preliminaries, followed by a detailed explanation of the ULTra framework for analyzing latent tokens.
+Additional details on the Transformer architecture are provided in Appendix G.
+3.1 Preliminaries
+Previous research on attention-based model interpretability [1, 8, 7, 52] has primarily focused on analyzing the semantic flow from input tokens to class logits.
+This is typically achieved by adding the attention probability matrix of each layer to an identity matrix and aggregating the results across attention heads through a weighted average.
+The weights are derived from the gradients of the class logits with respect to the attention probabilities.
+Specifically, to assess the contribution of each class logit using attention information, the contribution map for layer $b$ is defined as:
+$\mathbf{C}_{c}^{(b)}$ $=\mathbf{I}+\mathbb{E}_{h}\left[\left(\nabla_{\mathbf{A}_{h}^{(b)}}p(c)\right)^{+}\odot\mathbf{A}_{h}^{(b)}\right],$ (1)
+$\overline{S}_{c}$ $=\mathbf{C}_{c}^{(1)}\cdot\mathbf{C}_{c}^{(2)}\cdots\mathbf{C}_{c}^{(L)},$ (2)
+$S_{c}$ $=\overline{S}_{c}[0,1:],$ (3)
+where $\odot$ denotes the Hadamard product, and $(\cdot)^{+}$ is the operator that retains only positive values.
+$\mathbf{I}$ represent the identity matrix which reflect the contributions of skip connections and $\mathbf{A}_{h}^{(b)}\in\mathbb{R}^{(n+1)\times(n+1)}$ denotes the attention probability matrix of head $h$ and layer $b$.
+The term $\nabla_{\mathbf{A}^{(b)}_{h}}p(c)=\frac{\partial p(c)}{\partial\mathbf{A}_{h}^{(b)}}$ is the partial derivative of the attention map with respect to the predicted probability for class $c$, and $\mathbb{E}_{h}$ denotes the mean over multiple attention heads.
+Equation 2 represents matrix multiplication, which accounts for aggregating contributions from all layers.
+Each element of $\overline{S}_{c}[i,j]$ represents the influence of the $j$-th input token on the $i$-th output token.
+Thus, each element of $S_{c}[j]$ quantifies the influence of the $j$-th input token on class $c$, with token 0 corresponding to the CLS token.
+This raises a new question: Is there a similar approach to interpret latent tokens in Transformers?
+3.2 ULTra Framework
+Inspired by Equation 1, in this work we aim to measure the contribution of latent token $\mathbf{z}_{i}^{(l)}$ using the underlying attention probabilities, where $i\in\{0,1,\dots,n\}$ denotes the token index and $l$ represents the layer.
+Specifically, to quantify the influence of the attention map at layer $b$ on the latent token $\mathbf{z}_{i}^{(l)}$, the contribution map $\mathbf{C}_{i}^{(b,l)}$ is defined as follows:
+$\mathbf{C}_{i}^{(b,l)}=\mathbf{I}+\mathbb{E}_{h}\left(\left(\nabla_{\mathbf{A}^{b}_{h}}f({\mathbf{z}_{i}^{(l)}})\right)^{+}\odot\mathbf{A}^{b}_{h}\right),$ (4)
+Intuitively, this approach traces the most influential input information contributing to a latent token $\mathbf{z}_{i}^{(l)}$, where the notion of influence is defined based on the impact an input token has on the function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ of the target token.
+Specifically, by influence, we mean the degree to which an input token affects $f(\mathbf{z}_{i}^{(l)})$, reflecting its importance in the representation.
+Finally, we define the corresponding explanation map for the latent token $\mathbf{z}_{i}^{(l)}$, denoted as $S_{i}^{(l)}\in\mathbb{R}^{n}$, where each element represents the influence of an input token on $\mathbf{z}_{i}^{(l)}$.
+The explanation map is computed as:
+$\overline{S}_{i}^{(l)}$ $=\mathbf{C}_{i}^{(1,l)}\cdot\mathbf{C}_{i}^{(2,l)}\cdots\mathbf{C}_{i}^{(l-1,l)},$ (5)
+$S_{i}^{(l)}$ $=\overline{S}_{i}^{(l)}[i,1:],$ (6)
+where Equation 5 represents matrix multiplication, which aggregates contributions from all layers to the target latent token.
+Transformer skip connections cause most contributions to concentrate on $S_{i}^{(l)}[i-1]$, hindering token-level analysis.
+To mitigate this, we replace it with the maximum of the other elements, better capturing token contributions.
+Moreover, in some experiments, e.g., semantic segmentation, we reshape and upsample the explanation map using bilinear or cubic interpolation to match the input resolution, producing $\tilde{S}_{i}^{(l)}$.
+The overall framework of ULTra in shown in Figure 1.
+Additionally, As we illustrated in section 5, in our experimetns we utilize three approach for desiging $f$ in ULTra framework:
+(i) $f_{\mathbf{s}}({\mathbf{z}_{i}^{(l)}})=\langle\mathbf{z}_{i}^{(l)},\mathbf{1}\rangle$ In this formulation, $\mathbf{1}$ represents an all-ones vector, meaning that the function simply computes the sum of all elements in $\mathbf{z}_{i}^{(l)}$.
+This approach is denoted by $\text{ULTra}_{\scriptsize\mathcal{S}}\vphantom{p}$.
+The underlying intuition is that this approach treats all elements of the token equally.
+(ii) $f_{\mathbf{e}}({\mathbf{z}_{i}^{(l)}})=\langle\mathbf{z}_{i}^{(l)},\mathbf{z}_{i}^{(l)}\rangle$ This approach is based on the energy (or norm) of a given token, a concept that has been previously explored in the literature [16].
+By leveraging token energy, this method captures the magnitude of the token’s latent representation.
+We refer to this variant as $\text{ULTra}_{\scriptsize\mathcal{E}}\vphantom{p}$.
+(iii) $f_{\mathbf{w}}(\mathbf{z}_{i}^{(l)})=\langle\mathbf{z}_{i}^{(l)},\mathbf{w}_{i}\rangle$: Besides of previous approaches, we introduce a learnable vector $\mathbf{w}_{i}$ to project each token $\mathbf{z}_{i}^{(l)}$.
+Geometrically, the functionality of $f_{\mathbf{s}}$ in $\text{ULTra}_{\scriptsize\mathcal{S}}\vphantom{p}$ can be seen as mapping all tokens onto the same projection basis within the token space.
+Intuitively, the goal of this approach is to enhance the overall performance of the ULTra framework by computing a more informative projection basis.
+We refer to this variant as $\text{ULTra}_{\scriptsize\mathcal{W}}\vphantom{p}$.
+While $\text{ULTra}_{\scriptsize\mathcal{S}}\vphantom{p}$ is training-free, $\text{ULTra}_{\scriptsize\mathcal{W}}\vphantom{p}$ requires light training to learn the optimal $\mathbf{w}_{i}$ vectors.
+In our experiments, we train these vectors using a limited number of batches.
+Notably, we employ this approach to further enhance segmentation performance.
+Furthermore, in Section 4.1, we propose a self-supervised approach for learning $\mathbf{w}_{i}$.
+For both $\text{ULTra}_{\scriptsize\mathcal{S}}\vphantom{p}$ and $\text{ULTra}_{\scriptsize\mathcal{E}}\vphantom{p}$, our framework operates without requiring any additional training.
+In the case of semantic segmentation Experimental results demonstrate that these proposed approaches closely match—or even surpass—other baselines that rely on fine-tuning, highlighting their effectiveness and efficiency.
+4 Tailoring ULTra for Different Applications
+In this section, we aim to examine ULTra’s capability to adapt to various tasks involving semantic knowledge.
+Figure 2: ULTra segmentation results on sample images. The top row displays the original images, the middle row shows true annotations, and the bottom row presents our model’s predictions.
+(a) Original Image         (b) Latent Token’s explanation Map (c) Predicted Binary Mask
 Refer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
 (b) Overlay of $\tilde{S}_{i}^{(13)}$ on the original image for different $i$, where the location of the $i$-th token is indicated by the purple square.
-(c) The binary mask $M_{i}^{(13)}$ for each corresponding relevance map in (b).
-3 Methodology
-In this section, we present our approach for interpreting latent representations in Transformers.
-We begin with essential preliminaries and notation for clarity, followed by a detailed explanation of our method for interpreting latent tokens and its application to semantic segmentation.
-3.1 Preliminaries and Notation
+(c) The binary mask $M_{i}^{(13)}$ for each corresponding explanation map in (b).
+4.1 Unsupervised Semantic Segmentation
+Explanation maps are defined for each latent token at a fixed layer, resulting in a total of as many maps as there are latent tokens.
+For segmentation, we hierarchically cluster these maps.
+In our experiments, we set a predefined number of clusters $k$ for the algorithm.
+Additionally, in Section F, we conduct comprehensive experiments exploring the use of a threshold $\zeta$ to dynamically adjust the optimal cluster count $k$.
+To ensure fair representation, we apply min-max scaling to prevent larger objects from dominating the clustering process.
+After clustering, $k$ distinct concepts are defined by aggregating explanation maps. The aggregated explanation map for cluster $c$ is:
+$\tilde{S}^{(l)}_{c}[x,y]=\sum_{i\in\phi(c)}\tilde{S}_{i}^{(l)}[x,y],$ (7)
+where $\phi(c)=\{i:\text{Class}(\tilde{S}_{i}^{(l)})=c\}$ represents the grouping of label assignments.
+Given a fixed threshold $\zeta$, class labels are assigned to input pixels using the explanation map of the $l$-th layer as follows:
+$\text{Class}[x,y]^{(l)}=\underset{c\in\{1,\ldots,k\}}{\text{argmax}}\,S_{c}^{(l)}[x,y].$ (8)
+Some examples illustrating our segmentation method are presented in Figure 2.
+$\text{ULTra}_{\scriptsize\mathcal{W}}\vphantom{p}$.
+As previously mentioned, to further enhance segmentation performance, we propose $\text{ULTra}_{\scriptsize\mathcal{W}}\vphantom{p}$, which utilizes a learnable transformation matrix to obtain a more informative projection basis for the token space.
+For simplicity of notation, we fix the layer $\ell$ and omit the layer superscript in the following derivations.
+Consider $z_{i,j}=\mathbf{e}_{j}^{T}\mathbf{z}_{i}$, where $\mathbf{e}_{j}$ is the $j$-th standard basis vector.
+The gradients can then be rewritten as:
+$\nabla_{\mathbf{A}^{b}_{h}}f_{\mathbf{s}}(\mathbf{z}_{i})$ $=\sum_{j}\nabla_{\mathbf{A}^{b}_{h}}z_{i,j},$ (9)
+$\nabla_{\mathbf{A}^{b}_{h}}f_{\mathbf{w}}(\mathbf{z}_{i})$ $=\sum_{j}\mathbf{w}_{i,j}\nabla_{\mathbf{A}^{b}_{h}}z_{i,j}.$ (10)
+In other words, the gradient of $f_{\mathbf{w}}(\mathbf{z}_{i})$ is a weighted version of the gradient of $f_{\mathbf{s}}(\mathbf{z}_{i})$.
+From this perspective, by manipulating the weighting vector $\mathbf{w}_{i}$, we can prioritize the more informative dimensions of the token space.
+As illustrated in section 5, this leads to improved performance of ULTra compared to variants that rely on a naive choice of $f$.
+Here, a crucial question arises: How do we characterize the amount of information in the token space to achieve a good projection basis for the token space?
+In other words, we need to learn a matrix $\mathbf{W}=[\mathbf{w}_{1},\dots,\mathbf{w}_{n}]^{T}$ to perform more effective segmentation.
+To this end, we introduce a perturbation on the input image and, by optimizing $\mathbf{W}$, aim to minimize the divergence of specific tokens from their unperturbed versions.
+Let $\mathbf{x}$ be an input image and $\tilde{\mathbf{x}}$ its perturbed version corresponding to segmentation class $c$, which can be computed as:
+$\tilde{\mathbf{x}}^{c}=\mathbf{x}+\mathcal{P}_{\phi}(\mathbf{x},c;\delta),$ (11)
+where $\mathcal{P}_{\phi}$ [^1] The perturbation strategy is defined based on the model parameters $\phi$ since the model is involved in distinguishing the perturbation using its segmentation.
+is a perturbation strategy that applies perturbations to the corresponding negative regions predicted for class $c\in\mathcal{C}$.
+Moreover, $\delta$ quantifies the amount of perturbation, and for $\delta=0$, we have $\tilde{\mathbf{x}}^{c}=\mathbf{x}$.
+Now, to optimize the parameters $\mathbf{W}$, we formulate the problem as:
+$\min_{\mathbf{W}}$ $\mathbb{E}_{x\sim\mathcal{D}}\left[\sum_{c\in\mathcal{C}}\sum_{i\in\{j\mid\mathbf{x}_{j}=\mathbf{x}^{c}_{j}\}}d_{\mathbf{W}}(\mathbf{z}_{i},\tilde{\mathbf{z}}_{i})\right],$ (12)
+[…]
+where $d_{\mathbf{W}}(\mathbf{z}_{i},\tilde{\mathbf{z}}_{i})$ denotes the divergence between the two vectors based on the projection basis $\mathbf{w}_{i}$, $\|\cdot\|_{2}$ is the eculidian norm, and $\mathbf{x}_{i}$ and $\mathbf{z}_{i}$ represent the $i$-th input token and latent token, respectively.
+In particular, we do not perturb the regions recognized by $\text{ULTra}_{\scriptsize\mathcal{W}}\vphantom{p}$ that contain localized information.
+This encourages the model to become more aware of self-attended regions, ultimately enhancing its decision-making in these critical areas.
+In our experiments, we parameterize each projection vector as $\mathbf{w}_{k}=\boldsymbol{\theta}_{k}/\|\boldsymbol{\theta}_{k}\|_{2}$.
+Hence, based on the formulation in Equation 12, to optimize the parameters $\boldsymbol{\theta}_{k}$, we define the Self-Consistency loss as follows:
+$\mathcal{L}_{\text{sc}}(x,\boldsymbol{\Theta})=\sum_{c\in\mathcal{C}}\sum_{i\in\{j\mid\mathbf{x}_{j}=\mathbf{x}^{c}_{j}\}}\left|\mathbf{w}_{i}^{T}(\mathbf{z}_{i}-\tilde{\mathbf{z}}_{i})\right|^{2},$ (13)
+where $\mathbf{w}_{i}=\boldsymbol{\theta}_{i}/\|\boldsymbol{\theta}_{i}\|_{2}$ and $\mathbf{\Theta}=[\boldsymbol{\theta}_{1},\dots,\boldsymbol{\theta}_{n}]^{T}$.
+We use the term ”Self” because we leverage the model’s own segmentation results, obtained using $\text{ULTra}_{\scriptsize\mathcal{W}}\vphantom{p}$, to enhance its performance through a systematic weighting approach.
+Experimentally, we found that lightly optimizing $\mathbf{\Theta}$ on a limited number of samples improves the model’s segmentation performance.
+4.2 Latent Token Interpretability Assessment
+While semantic segmentation serves as an interpretability evaluation approach for ULTra, in this section, we evaluate ULTra in additional introduced metrics.
+Specifically, we designed two tasks: (i) Object Selection and (ii) Perturbation Test.
+Additionally, in Section 5.3, we propose quantitative metrics to evaluate our method on these tasks to provide deeper insights into the interpretability of ULTra by identifying influential regions and behaviors within the model’s decision-making process.
+Perturbation Test.
+The perturbation test is a widely used metric for evaluating explainability methods [8, 52].
+This approach involves perturbing parts of an image based on the explanation map generated by the method and measuring the resulting change in the predicted class probability.
+Inspired by this technique, instead of using class probability, we leverage the change in the embedding space to design a perturbation-based validation test.
+This allows us to assess the reliability of the explanation maps generated for the latent tokens.
+Object Selection.
+In this task, we convert the upsampled explanation map $\tilde{S}_{i}^{(l)}$ into a binary segmentation mask using a threshold $\tau$, where the binary mask $M_{i}^{(l)}$ is defined as:
+$M_{i}^{(l)}[x,y]=\begin{cases}0,&\text{if }\tilde{S}_{i}^{(l)}[x,y]<\tau,\\ 1,&\text{otherwise}.\end{cases}$ (14)
+Here, $\tilde{S}_{i}^{(l)}[x,y]$ represents the relevance value at position $[x,y]$ in $\tilde{S}_{i}^{(l)}$, with $\tau$ as the threshold.
+When $M_{i}^{(l)}[x,y]=1$, it indicates that the position $[x,y]$ belongs to the object region.
+Our findings indicate that as tokens propagate through the network, they refine their object representation while retaining the semantic meaning of their associated image patches performing Object Selection.
+Figure 4 visually illustrates this process.
+For a given patch token $\mathbf{x}_{i}$, the object it most strongly represents is denoted as class $k_{i}$.
+The latent token $\mathbf{z}_{i}^{(l)}$ generates a explanation map that highlights areas with higher values associated with class $k_{i}$ in the whole image, including $\mathbf{x}_{i}$.
+After applying a threshold, this map becomes a binary segmentation mask expected to exhibit a high Intersection over Union (IoU) with the corresponding class $k_{i}$ region in the image.
+An illustrative example is shown in Figure 3.
+(a) Original Image (b) Explanation Relevency Map
+Refer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+Figure 4: An example illustrating the model’s decision-making process across layers.
+Moving from left to right corresponds to deeper layers in the network.
+The first row corresponds to the CLS token, while the second, third, and fourth rows represent three different tokens, highlighted by red squares.
+4.3 Interpreting LLMs in Text Summarization
+In this section, we examine how our interpretability framework can be applied to text summarization tasks, taking steps toward uncovering the underlying intent of LLMs.
+In Section 5.4, we qualitatively evaluate $\text{ULTra}_{\scriptsize\mathcal{S}}\vphantom{p}$ using two samples by visualizing the regions of the input context that an LLM prioritizes while interpreting a given TL;DR summary.
+This analysis reveals key input regions shaping the model’s decisions, aiding in concise and relevant summary generation.
+For this task, we concatenate the context $x$ and the summary $y$ with a separator token.
+After feeding this input into the model, we compute the relevance scores of the TL;DR tokens with respect to the context tokens.
+We then average these scores for each token in $x$ to obtain a scalar value, referred to as the Token Contribution Score, $\lambda_{i}^{(l)}\in\mathbb{R}^{+}$, which highlights the contribution of each context token in interpreting the summary $y$ within the given context.
+Accordingly, $\lambda_{i}^{(l)}$ is computed as:
+$\lambda_{i}^{(l)}=\frac{1}{|y|}\sum_{j=1}^{|y|}S_{j+|x|}^{(l)}[i],\quad\forall i\in\{1,\cdots,|x|\},$ (15)
+where ||$\cdot$|| ⋅ | denotes the number of tokens in the text.
+5 Experiments & Results
+5.1 Experimental Setup
+Datasets.
+In our experiments, we evaluate model performance on several semantic segmentation benchmarks, focusing on vision-related tasks.
+We conducted experiments on four datasets: COCO-Stuff 27 [5], PASCAL VOC 2012 [19], Potsdam-3 [21], and Cityscapes [14].
+For our qualitative analysis of LLM interpretation in the task of text summarization, as described in Section 4.3, we utilized the TL;DR dataset [44].
+The TL;DR dataset contains summary comparisons with human feedback collected by OpenAI.
+Models.
+For all experiments in the vision tasks, we used different pretrained versions of CLIP’s image encoder [36] as well as DINO ViT-S/16 and ViT-B/16 [6].
+For interpreting text summarization, as described in Section 4.3, we used the Llama-2-7B language model [47].
+All experiments were run on 4 NVIDIA A100-80GB GPUs.
+Further details of the datasets and models used are provided in Appendix C.
+Method Model Training U. ACC U. mIoU
+IIC [23] R18+FPN 21.8 6.7
+PiCIE [11] R18+FPN 48.1 13.8
+DINO [6] ViT-S/8 28.7 11.3
+ViT-S/16 22.0 8.0
+ViT-B/8 30.5 9.6
+ACSeg [27] ViT-S/16 - 16.4
+TransFGU [55] ViT-S/8 52.7 17.5
+STEGO [20] ViT-S/8 48.3 24.5
+ViT-S/16 52.5 23.7
+ViT-B/8 56.9 28.2
+STEGO +HP [41] ViT-S/8 57.2 24.6
+ViT-S/16 54.5 24.3
+DepthG [42] ViT-S/8 56.3 25.6
+ViT-B/8 58.6 29.0
+U2Seg [32] R50 63.9 30.2
+ViT-B/32 60.6 $\uparrow$2 34.1 $\uparrow$5.1
+$\text{ULTra}_{\mathcal{W}}^{\text{CLIP}}\vphantom{p}$ ViT-B/16 63.8 $\uparrow$5.2 34.0 $\uparrow$5
+ViT-L/14 67.9 38.2
+ViT-B/32 60.8 $\uparrow$2.2 34.6 $\uparrow$5.6
+$\text{ULTra}_{\mathcal{S}}^{\text{CLIP}}\vphantom{p}$ ViT-B/16 63.0 $\uparrow$4.4 33.2 $\uparrow$4.2
+ViT-L/14 66.5 37.5
+ViT-S/16 67.2 $\uparrow$10.9 34.4 $\uparrow$8.8
+$\text{ULTra}_{\mathcal{W}}^{\text{DINO}}\vphantom{p}$ ViT-B/16 67.4 $\uparrow$8.8 37.7 $\uparrow$8.7
+ViT-S/16 66.4 $\uparrow$7.8 33.3 $\uparrow$7.7
+$\text{ULTra}_{\mathcal{S}}^{\text{DINO}}\vphantom{p}$ ViT-B/16 67.3 $\uparrow$8.7 35.6 $\uparrow$6.6
+Table 1: Comparison of different unsupervised segmentation methods on the COCO-Stuff dataset.
+Method Model Training U. mIoU
+IIC [23] R18+FPN 9.8
+MaskContrast [48] R50 35.0
+Leopart [57] ViT-S/16 41.7
+TransFGU [55] ViT-S/8 37.2
+MaskDistill [49] ViT-S/16 + R50 42.0
+ACSeg [27] ViT-S/16 47.1
+ViT-B/32 51.2
+$\text{ULTra}_{\mathcal{W}}^{\text{CLIP}}\vphantom{p}$ ViT-B/16 50.9
+ViT-L/14 49.1
+ViT-B/32 49.2
+$\text{ULTra}_{\mathcal{S}}^{\text{CLIP}}\vphantom{p}$ ViT-B/16 48.3
+ViT-L/14 48.7
+ViT-S/16 48.9 $\uparrow$1.8
+$\text{ULTra}_{\mathcal{W}}^{\text{DINO}}\vphantom{p}$ ViT-B/16 50.5
+ViT-S/16 47.9 $\uparrow$0.1
+$\text{ULTra}_{\mathcal{S}}^{\text{DINO}}\vphantom{p}$ ViT-B/16 50.0
+Table 2: Comparison of different unsupervised segmentation methods on the PASCAL VOC 2012 dataset.
+5.2 Semantic Segmentation
+We benchmarked the segmentation performance of our method against several SOTA approaches in the literature on unsupervised segmentation.
+To train the matrix $W$ for $\text{ULTra}_{\scriptsize\mathcal{W}}\vphantom{p}$, we utilized 256 samples.
+To assess the effectiveness of our method, we employed the Unsupervised mean Intersection over Union (U.
+mIoU) and Unsupervised pixel Accuracy (U.
+ACC) metrics.
+Our experimental results on the COCO-Stuff and PASCAL VOC datasets for $\text{ULTra}_{\scriptsize\mathcal{W}}\vphantom{p}$ and $\text{ULTra}_{\scriptsize\mathcal{S}}\vphantom{p}$ are presented in Tables 1 and 2.
+In these tables, the ”Training” column indicates whether additional training is required, and the up/down arrows denote the performance difference between our method and the state-of-the-art for the same model.
+[^2] To ensure a fair comparison based on model architecture, we consider only the number of model parameters.
+For instance, all ViT-B models with different patch sizes are compared against each other.
+Additionally, for some models, such as ViT-L/14, no existing baseline is available for direct comparison, highlighting the versatility of ULTra across different architectures.
+Furthermore, the comparative results for the Cityscapes and Potsdam datasets are provided in Appendix D.1.
+The results of $\text{ULTra}_{\scriptsize\mathcal{E}}\vphantom{p}$ are also presented in Appendix D.3.
+Moreover, we conducted an ablation study related to the model depth in Appendix A.
+5.3 Interpretability Evaluations
+Model Perturb. PASCAL VOC COCO-Stuff
+Neg. $\downarrow$ Pos. $\uparrow$ Neg. $\downarrow$ Pos. $\uparrow$
+ViT-B/32 Mask 13.16 15.51 12.62 15.21
+Noise 6.97 10.23 6.83 10.26
+ViT-B/16 Mask 14.68 17.5 14.16 17.39
+Noise 9.87 14.03 10.14 14.12
+ViT-L/14 Mask 6.66 9.43 6.38 9.31
+Noise 4.45 7.88 4.27 7.74
+Table 3: Average token vector differences for ViT models under mask and noise perturbation tests on PASCAL VOC and COCO-Stuff datasets, highlighting the impact of positive and negative perturbations based on the relevancy map on token representations.
+Refer to caption Refer to caption
+(a) TL;DR: I’ve had very bad luck with girls my whole (b) TL;DR: I still like her but my rational
+life and I don’t know how to get my confidence up. side says ”no, she is a trash person”.
+Figure 5: Visualization of Token Contribution Scores ($\lambda_{i}^{(l)}$) highlighting the relevance of context tokens in interpreting the summary.
+Each token is colored proportionally to its $\lambda_{i}^{(l)}$ value.
+These visualizations demonstrate the model’s ability to identify key semantic elements in the context for generating relevant summaries.
+Perturbation Test.
+To assess the reliability of the explanation maps, we conduct a perturbation test by selectively altering image regions based on the explanation map $S_{i}^{(l)}$ for each token $i$.
+The perturbation is applied at the patch level while ensuring the total perturbed area remains consistent across all cases.
+We consider two types of perturbations:
+(i) Positive Perturbation: Removing highly relevant regions, which should significantly affect the token’s representation.
+(ii) Negative Perturbation: Removing less relevant regions, which should have minimal impact.
+To ensure fairness, the initial token’s patch remains unchanged in both perturbations, as its information is directly propagated to the target token through skip connections.
+Given a perturbed representation $\tilde{z}_{i}^{(l)}$, we measure the deviation from the original token representation $z_{i}^{(l)}$ using the Euclidean distance, then we compute the average deviation over the entire dataset by selecting k random tokens per image and aggregating the distances:
+$\bar{d}_{\text{Euc}}=\frac{1}{M}\sum_{j=1}^{M}\left(\frac{1}{k}\sum_{i=1}^{k}\|z_{i,j}^{(l)}-\tilde{z}_{i,j}^{(l)}\|_{2}\right),$ (16)
+where $M$ is the number of images, and $k$ tokens are randomly selected from each.
+For our results, we set $k=10$.
+The perturbation results for COCO-Stuff and PASCAL VOC are presented in Table 3, While a full table covering all datasets can be found in Appendix D.2.
+Positive perturbations result in greater deviations compared to negative ones, confirming that the explanation maps effectively highlight influential regions.
+The consistency of these results across multiple datasets further validates the reliability of our interpretability framework.
+Object Selection.
+To quantify alignment, we compute the IoU by converting the explanation map $S_{i}^{(l)}$ into a binary mask $M_{i}^{(l)}$ and comparing it with the ground-truth mask.
+We propose the Initial Token IoU (ITIoU) metric, which measures how well the explanation maps of input tokens align with their respective class masks.
+The ITIoU is calculated as:
+$\textit{ITIoU}^{(l)}(X)=\frac{1}{C}\sum_{i=1}^{C}\frac{1}{|\mathcal{T}_{i}|}\sum_{\mathbf{x}_{j}\in\mathcal{T}_{i}}\text{IoU}(M_{j}^{(l)},G_{i}),$ (17)
+where $C$ denotes the number of classes, $\mathcal{T}_{i}$ represents the set of tokens associated with class $i$, $M_{j}^{(l)}$ is the binary segmentation mask for token $\mathbf{x}_{j}$ within class $i$, and $G_{i}$ is the ground-truth mask for class $i$ in image $\mathbf{x}$.
+The inner sum averages the IoU for tokens in $\mathcal{T}_{i}$ for each class, and the outer sum then averages across all classes.
+Using a threshold of 0.2, our ITIoU [^3] A more detailed analysis of ITIoU is provided in Appendix E.
+metric achieves an average score of 37.84 % on the COCO-Stuff validation dataset and 39.51 % on the PASCAL VOC dataset.
+5.4 Interpretable Text Summarization
+In this experiment, we used a Supervised Fine-Tuned (SFT) version of Llama-2-7B trained on the UltraFeedback Binarized (UFB) dataset [15].
+Additionally, we aligned the model to the text summarization task on the TL;DR dataset using the Direct Preference Optimization (DPO) method [37] for 1,000 iterations, with a learning rate of $5\times 10^{-6}$ and $\beta=0.5$.
+To validate our framework, we selected the preferred response (TL;DR) of each sample in the dataset, denoted by $y$, and used it as the summary of the context $x$.
+In example (a): semantically significant words such as ‘relationship’, ‘experience’, ‘rejection’, and ‘never’ are prominently highlighted, reflecting the model’s interpretation of the person’s struggles with relationships and feelings of rejection.
+Additionally, the highlighting of the question at the beginning of the context ‘How do I stop feeling bad…’ suggests the model recognizes the presence of uncertainty and a request for guidance, which is encapsulated in the summary as ‘I don’t know.’
+In example (b): $\lambda_{i}^{(l)}$ scores reveals the model’s focus on words such as ‘feelings’, ‘hate’, ‘disappoint’, ‘love’, and ‘like’, which correspond to the person’s mixed emotions toward their girlfriend, as described in the summary.
+The apparent contradiction between ‘love’ and ‘trashness’ in the summary seems to be derived from these highlighted terms, suggesting the model understands the conflicting emotions present in the text.
+.
+Furthermore, the focus on ‘character’ reflects the summary’s judgmental tone, implying the model links it to personality assessment.
+This token-level analysis visually illustrates the model’s context processing, aiding interpretable fine-tuning and alignment methods like RLHF [13, 44, 33] and Preference Optimization [37, 2, 18], where understanding model behavior and intent is essential.
+6 Concluding Remarks
+Summary.
+We present a framework for interpreting latent tokens in Transformers, achieving SOTA performance in unsupervised semantic segmentation.
+Additionally, we validate our approach using perturbation tests and object selection, demonstrating its broader applicability.
+Future Research.
+As a work on latent token interpretability, our framework extends existing methods.
+Future directions include refining it for latent tokens and addressing its computational complexity at inference time (Appendix B.2).
+Appendix A Study on the Effect of Layer Depth in ViT Token Understanding
+In this section, we analyze the impact of depth on our model’s interpretability and segmentation performance, providing insights into the contribution of each layer.
+As anticipated and observed in Figure 7, deeper layers generally carry more semantic significance.
+However, the contribution diminishes in the final layers, suggesting that a depth of around 13 layers might be more than sufficient for the ViT to effectively comprehend image content.
+This finding implies that even fewer layers might achieve comparable results, potentially reducing computational costs without compromising performance.
+We observe an intriguing behavior in the initial layers, where performance initially declines before improving.
+This phenomenon is also visually evident in Figure 4, where the attention maps in the first layer appear to focus on the entire image.
+This suggests that, initially, the token examines the image as a whole before selectively gathering information from tokens with similar characteristics.
+In the CLIP ViT-L/14 model, deeper layers capture more refined and detailed feature representations.
+Unlike shallower models, where segmentation performance stabilizes early, ViT-L/14 benefits from its depth by gradually extracting richer hierarchical features.
+As shown in Figure 6, accuracy and Mean IoU improve as layers deepen, with segmentation performance peaking around the mid-depth layers.
+However, in deeper layers (beyond layer 20), segmentation quality slightly declines.
+This suggests that while the model gains a better understanding of high-level semantics, it loses some spatial precision.
+This trade-off occurs because later layers prioritize semantic abstraction over detailed spatial structures.
+While deeper models like ViT-L/14 offer improved feature extraction, more layers do not always lead to better segmentation.
+Instead, an optimal balance between depth and spatial representation is necessary for effective segmentation.
+Figure 6: The segmentation performance of the CLIP ViT-L/14 model in an unsupervised setting, measured by accuracy and IoU, changes across different layers.
+As layer depth increases, both metrics improve at first, reaching their highest point in the mid-depth layers.
+However, performance slightly declines in deeper layers.
+This suggests that while deeper layers capture richer semantic details, they also introduce complexity that does not always improve segmentation.
+Additionally, the limitations of labeled datasets, especially the ambiguity in object definitions, further restrict the model’s ability to achieve better segmentation, despite its strong interpretability.
+Refer to caption Refer to caption
+(a) COCO-Stuff (b) PASCAL VOC
+Figure 7: Ablation study on two evaluation metrics across layers.
+These plots demonstrate a progressive improvement in semantic segmentation performance in the deeper layers of the transformer model.
+This enhancement is attributed to latent tokens capturing more meaningful segment structures, resulting in increasingly accurate and refined semantic representations.
+Appendix B Limitations & Discussions
+B.1 Datasets
+A significant challenge in deep learning is providing labeled datasets, which are essential for the success of deep learning methods.
+However, labeling can be ambiguous, as objects or attributes may be labeled separately or combined as a single entity.
+Our method demonstrates how ViT interprets images in a zero-shot setting, capturing fine-grained or general representations.
+Notably, our method’s predictions often exhibit logical consistency that surpasses the ground truth, effectively identifying relationships between objects.
+However, since our approach does not involve supervised training, it cannot adapt itself to detect only the specific objects labeled in the dataset.
+As a result, while the logical quality of the predictions is high, the numerical metrics may decline due to mismatches with the dataset’s ground truth labels.
+Figure 8 illustrates this phenomenon with examples, showing instances where our method successfully captures unlabeled objects that are omitted in the dataset annotations.
+Figure 8: This figure illustrates challenging samples of datasets COCO-Stuff 27, Potsdam-3, and PASCAL VOC that contribute to low performance due to poor dataset labeling quality.
+The first row displays the original images, the second row shows the ground truth labels and the third row presents our predicted segmentation.
+B.2 Method
+In this work, we aimed to uncover semantic information embedded within the latent tokens of transformers.
+A key challenge in interpretability is the evaluation process.
+In traditional model interpretation, where only class logits are available as the primary source of information, various evaluation approaches have been proposed.
+However, in our case, many of these metrics are not directly applicable due to the fundamental differences in our methodology.
+To address this, we evaluated our approach in the context of unsupervised semantic segmentation.
+Although our baseline methods were originally designed for this task, we achieved superior results by aggregating our explanation maps.
+This outcome highlights the potential for future research in this direction.
+Specifically, improved aggregation methods—such as clustering techniques—could enhance performance, particularly when dealing with smaller patch sizes.
+Notably, in the case of a patch size of 8, clustering algorithms failed due to the high number of tokens, making effective aggregation more challenging.
+Moreover, despite being a zero-shot method, our approach incurs a significant computational cost when applied to semantic segmentation.
+This is primarily due to the necessity of computing multiple gradients for generating explanation maps, a process that must be performed for each token.
+As a result, inference time is considerably high.
+Future work could explore more efficient strategies to mitigate this issue, such as leveraging approximation techniques or optimizing the gradient computation process.
+Appendix C Datasets & Models
+C.1 Datasets
+We utilize a combination of datasets to provide a diverse testing ground for evaluating our method across both standard and challenging perspectives in semantic segmentation and interpretability evaluations.
+  • COCO-Stuff 27 [29]: A subset of the COCO dataset, featuring complex real-world scenes with pixel-level annotations across various object categories.
+  • PASCAL VOC 2012 [19]: A widely used benchmark containing pixel-level annotations for foreground objects in structured scenes.
+  • Potsdam-3: A high-resolution aerial-view dataset capturing urban landscapes, including buildings, roads, and vegetation, presenting additional challenges due to its large-scale top-down perspective.
+  • Cityscapes [14]: An urban street scene dataset with fine-grained pixel-level annotations, enabling the evaluation of segmentation performance in structured environments.
+These datasets allow for a comprehensive evaluation of our approach across different environments, ensuring robustness across diverse segmentation challenges.
+C.2 Models
+We employ various Vision Transformers (ViTs) pre-trained on large-scale datasets.
+These models process images as non-overlapping patches and employ self-attention mechanisms across multiple layers to capture long-range dependencies.
+Transformer Architectures
+  • CLIP ViT [36]: A vision transformer trained using contrastive learning on 400 million image-text pairs. It encodes images into a shared embedding space with text prompts. CLIP variants include:
+    • ViT-B/16: Consists of 12 transformer layers, a hidden size of 768, and processes images with 16×16 patch resolution.
+    • ViT-B/32: Similar to ViT-B/16 but with a 32×32 patch resolution, reducing computational cost at the expense of finer details.
+    • ViT-L/14: A larger model with 24 transformer layers, a hidden size of 1024, and a 14×14 patch resolution, providing enhanced feature extraction.
+  • DINO ViT [6]: A self-supervised vision transformer trained using knowledge distillation without labeled data.
+It learns image representations by maximizing similarity between different augmented views.
+Evaluated variants:
+    • ViT-S/16: A smaller model with 12 transformer layers, a hidden size of 384, and a patch size of 16×16.
+    • ViT-B/16: A larger model with 12 layers, a hidden size of 768, and a patch size of 16×16, providing stronger feature representation.
+These models serve as a robust testbed for evaluating both semantic segmentation and interpretability, ensuring our method generalizes across multiple architectures and dataset distributions.
+Appendix D Further Experiments
+D.1 Results of Semantic Segmentation on Cityscapes and Potsdam Datasets
+Tables 5 and 5 present the performance of ULTra on the Cityscapes and Potsdam datasets.
+Our method achieves strong segmentation performance, consistently outperforming prior unsupervised approaches on both datasets.
+Notably, the results demonstrate that ULTra effectively captures the semantic structure of the input data, achieving strong performance in semantic segmentation.
+Method Model Training U. mIoU
+IIC [23] R18+FPN 6.4
+PiCIE [11] R18+FPN 12.3
+STEGO [20] ViT-B/8 21.0
+STEGO +HP [41] ViT-S/8 18.4
+ViT-B/8 18.4
+DepthG [42] ViT-B/8 23.1
+ViT-B/32 17.6 $\downarrow$5.5
+$\text{ULTra}_{\mathcal{W}}^{\text{CLIP}}\vphantom{p}$ ViT-B/16 24.8 $\uparrow$1.7
+ViT-L/14 25.1
+ViT-B/32 17.1 $\downarrow$6
+$\text{ULTra}_{\mathcal{S}}^{\text{CLIP}}\vphantom{p}$ ViT-B/16 24.2 $\uparrow$1.1
+ViT-L/14 24.2
+ViT-S/16 25.8 $\uparrow$7.4
+$\text{ULTra}_{\mathcal{W}}^{\text{Dino}}\vphantom{p}$ ViT-B/16 26.5 $\uparrow$3.4
+ViT-S/16 24.2 $\uparrow$5.8
+$\text{ULTra}_{\mathcal{S}}^{\text{Dino}}\vphantom{p}$ ViT-B/16 25.7 $\uparrow$2.6
+Table 4: Comparison of different unsupervised segmentation methods on the Cityscapes dataset.
+Method Model Training U. ACC
+IIC [23] R18+FPN 65.1
+DINO [6] ViT-S/8 71.3
+STEGO [20] ViT-S/8 77.0
+DepthG [42] ViT-S/8 80.4
+ViT-B/32 78.7
+$\text{ULTra}_{\mathcal{W}}^{\text{CLIP}}\vphantom{p}$ ViT-B/16 80.9
+ViT-L/14 82.4
+ViT-B/32 78.3
+$\text{ULTra}_{\mathcal{S}}^{\text{CLIP}}\vphantom{p}$ ViT-B/16 80.9
+ViT-L/14 82.8
+ViT-S/16 79.0 $\downarrow$1.4
+$\text{ULTra}_{\mathcal{W}}^{\text{Dino}}\vphantom{p}$ ViT-B/16 80.7
+ViT-S/16 77.4 $\downarrow$3
+$\text{ULTra}_{\mathcal{S}}^{\text{Dino}}\vphantom{p}$ ViT-B/16 80.8
+Table 5: Comparison of different unsupervised segmentation methods on the Potsdam dataset.
+D.2 Complete Results of Perturbation Test
+We analyze how perturbations affect token representations by modifying patches at the input level.
+Figure 9 provides an example from the PASCAL VOC dataset, where the model used is CLIP ViT-B/32.
+The perturbations are applied consistently across all models while ensuring that the initial token remains unchanged, preserving global feature encoding.
+Two types of modifications are introduced.
+In the masking perturbation, selected patches are replaced with zeros, effectively removing the corresponding visual information.
+This is applied to both highly relevant regions (positive masking) and less relevant areas (negative masking).
+In contrast, the noise perturbation introduces Gaussian noise to the same sets of patches, adding controlled randomness to test the robustness of token representations.
+The noise follows a standard normal distribution with a perturbation strength of 0.3.
+To evaluate the impact of these perturbations, we compute the Euclidean distance between original and perturbed token representations.
+(a) Original
+(b) Positive Mask
+(c) Positive Noise
+(d) explanation Map
+(e) Negative Mask
+(f) Negative Noise
+Figure 9: Effect of perturbations on a sample image.
+(a) Original image, (b) Positive mask (removing key regions), (c) Positive noise, (d) explanation map, (e) Negative mask (removing less relevant areas), (f) Negative noise.
+Table 6 presents the results of $\text{ULTra}_{\scriptsize\mathcal{S}}\vphantom{p}$ on the Perturbation Test across four datasets.
+Model Perturb. VOC Potsdam COCO Cityscapes
+Neg. $\downarrow$ Pos. $\uparrow$ Neg. $\downarrow$ Pos. $\uparrow$ Neg. $\downarrow$ Pos. $\uparrow$ Neg. $\downarrow$ Pos. $\uparrow$
+ViT-B/32 Mask 13.16 15.51 12.35 14.02 12.62 15.21 9.95 14.42
+Noise 6.97 10.23 9.39 11.44 6.83 10.26 5.99 11.13
+ViT-B/16 Mask 14.68 17.5 13.91 15.51 14.16 17.39 11.33 16.27
+Noise 9.87 14.03 13.5 14.45 10.14 14.12 9.46 14.22
+ViT-L/14 Mask 6.66 9.43 6.61 8.99 6.38 9.31 5.36 8.86
+Noise 4.45 7.88 4.95 7.96 4.27 7.74 3.89 7.77
+Table 6: Average token vector differences for ViT models under mask and noise perturbation tests across multiple datasets, highlighting the impact of positive and negative perturbations based on the relevancy map on token representations.
+D.3 Result of $\text{ULTra}_{\mathcal{W}}$ Variant
+In this section, we present the performance of $\text{ULTra}_{\scriptsize\mathcal{E}}\vphantom{p}$ across COCO-Stuff, PASCAL VOC, Potsdam, and Cityscapes.
+While the results of this variant are lower than $\text{ULTra}_{\scriptsize\mathcal{S}}\vphantom{p}$, it still achieves competitive performance compared to the baseline methods.
+The detailed results are shown in Table 7.
+Method Model COCO VOC CityScapes Potsdam
+U. ACC U. mIoU U. mIoU U. mIoU U. ACC
+[…]
+ViT-B/16 53.6 26.6 40.0 20.9 70.4
+ViT-L/14 59.0 31.7 45.2 23.0 75.5
+[…]
+ViT-B/16 63.0 31.3 46.7 23.0 76.8
+Table 7: Results of ULTraE across different datasets.
+Appendix E ITIoU Analysis
+This section evaluates the effectiveness of ITIoU in assessing the performance of our object selection process.
+As expected, the final layers exhibit superior performance compared to the earlier layers, consistent with the results illustrated in Figure 11.
+This improvement highlights the increasing relevance of features in deeper layers for accurate object selection.
+Additionally, we observe the existence of an optimal threshold, $\tau$, which significantly influences the segmentation performance.
+This phenomenon is depicted in Figure 11, where performance trends are analyzed across different threshold values.
+Figure 10: Impact of varying the threshold $\tau$ on ITIoU performance for COCO-Stuff 27 dataset. The plot demonstrates the existence of an optimal $\tau$, where segmentation performance is maximized.
+Figure 11: Layer-wise ITIoU analysis for COCO-Stuff 27 dataset.
+Final layers perform significantly better than earlier ones due to their ability to capture high-level semantic features.
+This progression is evident in the increasing ITIoU values.
+Appendix F Threshold-Based Segmentation
+Dataset U. ACC U. mIoU
+COCO-Stuff 27 67.2 32.9
+PASCAL VOC - 51.9
+Potsdam 74.6 -
+Table 8: ULTra results across different datasets using the same threshold $\zeta=0.4$. Accuracy (U. ACC) and Mean IoU (U. mIoU) are reported where available.
+Hierarchical clustering is used to segment the token explanation maps, where the threshold parameter $\zeta$ controls the level of granularity.
+Lower values of $\zeta$ produce fine-grained segmentations, while higher values merge similar clusters, leading to broader groupings.
+Figure 13 illustrates the hierarchical clustering tree, where each leaf node represents a single token explanation map.
+As $\zeta$ increases, multiple explanation maps are grouped into larger clusters, reducing segmentation granularity.
+The rightmost part of the figure shows how different $\zeta$ values affect the final segmentation output.
+Figure 12 further quantifies the impact of $\zeta$ on segmentation performance for the COCO-Stuff 27 dataset.
+Accuracy and Mean IoU metrics are plotted as a function of $\zeta$, demonstrating that while different values of $\zeta$ yield varying levels of segmentation detail, our method remains robust across a range of threshold values.
+Figure 12: Impact of the clustering threshold $\zeta$ on segmentation performance for the COCO-Stuff 27 dataset.
+Accuracy and Mean IoU metrics show how different values of $\zeta$ affect segmentation quality.
+The results confirm that selecting an appropriate $\zeta$ enables the model to segment objects at different levels of abstraction, demonstrating the adaptability of pre-trained Vision Transformers in unsupervised semantic segmentation.
+Table 8 summarizes the segmentation performance of our method across multiple datasets using a fixed threshold $\zeta=0.4$.
+However, as discussed in Section B.1, not all threshold values may be equally suited for standard datasets.
+${\zeta=0.5}$${\zeta=0.35}$${\zeta=0.2}$
+Segmentation Results
+Original Image
+Figure 13: Hierarchical clustering tree showing the grouping of token explanation maps for all tokens in a latent layer of the Vision Transformer, not limited to the CLS token.
+Each leaf node represents a single token explanation map, while higher-level nodes show aggregated clusters based on a clustering threshold $(\zeta)$, which controls the level of detail.
+Lower $\zeta$ values reveal finer details, while higher values create broader, more general clusters.
+Appendix G Transformer Architecture
 The architecture of a typical Transformer can be formulated as follows: the input $X$ is split into $n$ tokens $\{\mathbf{x}_{i}\}_{i=1}^{n}$.
@@
 Here, $W_{1}^{(l)}\in\mathbb{R}^{d\times m}$, $W_{2}^{(l)}\in\mathbb{R}^{m\times d}$.
-3.2 Interpreting Latent Tokens
-A straightforward approach for interpreting a model involves analyzing the semantic flow from the input to the corresponding logits.
-This can be achieved by adding the attention probability matrix to an identity matrix $I$, which incorporates skip connections, and then multiplying the result across layers [1].
-However, a notable challenge arises from multiple attention heads in each Transformer layer.
-To address this, [6] proposes performing a weighted average across attention heads, with the weights determined by the gradient of the logits with respect to the attention weights.
-We employed a modified, simpler version of this method, utilizing only the attention weights rather than the layer’s relevance.
-In our framework, latent tokens $\mathbf{z}_{i}^{(l)}$ are represented as vectors rather than direct class correspondences, introducing an additional layer of complexity.
-To manage this, an appropriate transformation, such as the energy or euclidean norm of the latent token, is employed as a surrogate.
-Consequently, we compute the gradient of $\|\mathbf{z}_{i}^{(l)}\|$ with respect to the attention weights to facilitate the analysis.
-In the following equations, we define the relevance map $S_{i}^{(l)}\in\mathbb{R}^{n}$, where the $j$-th element of this vector represents the importance of the $j$-th input token to the targeted latent token $\mathbf{z}_{i}^{(l)}$.
-The relevance map is computed as:
-$\overline{\mathbf{A}}_{i}^{(b,l)}$ $=I+\mathbb{E}_{h}\left(\nabla_{\mathbf{A}^{b}_{h}}\lVert\mathbf{z}_{i}^{(l)}\rVert\odot\mathbf{A}^{(b)}_{h}\right)^{+},$
-$\overline{S}_{i}^{(l)}$ $=\overline{\mathbf{A}}_{i}^{(1,l)}\cdot\overline{\mathbf{A}}_{i}^{(2,l)}\cdot\;\dots\;\cdot\overline{\mathbf{A}}_{i}^{(l-1,l)},$
-$S_{i}^{(l)}$ $=\overline{S}_{i}^{(l)}[i,1:\;],$ (7)
-where $(\cdot)^{+}$ means considering only positive values, $S,\;\overline{\mathbf{A}}_{i}^{(b,l)}\in\mathbb{R}^{(n+1)\times(n+1)}$, $\odot$ denotes the Hadamard product, $\mathbb{E}_{h}$ represents the mean across the heads dimension, and $\mathbf{A}^{(b)}_{h}\in\mathbb{R}^{h\times(n+1)\times(n+1)}$ is the attention score matrix in the $b$-th layer and $A_{h,i,j}^{(b)}=\alpha_{h,i,j}^{(b)}$.
-Due to the skip connections in the transformer, most of the contribution of $S_{i}^{(l)}$ is concentrated on $S_{i}^{(l)}[i-1]$ which makes it hard to analyze other tokens’ contribution.
-To address this issue, we replace this element with the maximum value of other elements, thereby capturing the contributions of additional tokens to the selected token.
-For vision tasks, we first reshape the relevance map and then upsample it using bilinear or cubic interpolation to match the resolution of the model’s input.
-The resulting higher-dimensional matrix is denoted as $\tilde{S}$.
-This upsampling step is essential for enabling accurate object selection and semantic segmentation tasks.
-3.3 ULTra in Unsupervised Tasks
-In this section, we aim to examine ULTra’s capability to adapt to various tasks involving semantic knowledge.
-Importantly, it requires no additional training, leveraging the inherent understanding within transformers rather than relying on loss functions objective, final layer outputs, or fine-tuning.
-Unsupervised Semantic Segmentation.
-As previously discussed, a relevance map can be defined for each latent token at a fixed layer, with the total number of relevance maps equal to the number of latent tokens.
-In the context of segmentation, the goal is to assign a class label to each pixel within an image.
-To achieve this, we employ clustering techniques, such as hierarchical clustering, that do not require a predefined number of classes.
-These techniques group the relevance maps into $k$ distinct clusters, where $k$ is unknown.
-Ideally, we aim for $k$ to approximate the actual number of classes present in the image.
-Our approach provides flexibility in adjusting the value of $k$ by modifying the cutoff distance threshold $\zeta$ within the clustering algorithm.
-Increasing $\zeta$ produces fewer, broader clusters that capture general categories, such as background and foreground.
-Conversely, reducing $\zeta$ allows for finer segmentation, distinguishing more specific features, such as an object’s head or hands.
-To prevent the method from disproportionately favoring larger objects, given that the number of elements in each cluster may vary, we apply min-max scaling to each cluster independently.
-After clustering, we define $k$ distinct concepts by aggregating the relevance maps within each cluster. The aggregated relevance map $\tilde{S}^{(l)}_{c}[x,y]$ for a cluster $c$ is computed as:
-$\tilde{S}^{(l)}_{c,\zeta}[x,y]=\sum_{i\in\phi_{\zeta}(c)}\tilde{S}_{i}^{(l)}[x,y],$ (8)
-where $\phi_{\zeta}(c)=\{i\mid\text{Class}(\tilde{S}_{i}^{(l)})=c\}$ represents the set of label assignments determined by the clustering algorithm.
-For each pixel at position $[x,y]$, the class label is determined by identifying the cluster $c$ with the highest relevance value at that pixel.
-Mathematically, the class assignment for a pixel is expressed as:
-$\text{Class}[x,y]_{\zeta}^{(l)}=\underset{c\in\{1,\ldots,k\}}{\text{argmax}}\,S_{c,\zeta}^{(l)}[x,y].$ (9)
-Some examples illustrating our segmentation method is presented in Figure 3. Additionally, the hierarchy tree and the effect of the threshold are demonstrated in Figure 1.
-Figure 3: ULTra segmentation results on sample images. The top row displays the original images, the middle row shows ground-truth annotations, and the bottom row presents our model’s predictions.
-4 Experiment
-Datasets.
-In our experiments, we evaluate model performance on several semantic segmentation benchmarks, focusing on vision-related tasks.
-We conducted experiments on three datasets: COCO-Stuff 27 [4], PASCAL VOC 2012 [15], and Potsdam-3 [19].
-This combination of datasets provides a diverse testing ground to rigorously evaluate our unsupervised zero-shot approach across both standard and challenging perspectives in semantic segmentation.
-COCO-Stuff 27, a subset of the COCO dataset [28], includes complex, real-world scenes with pixel-level annotations across various object categories.
-Similarly, PASCAL VOC 2012 serves as a classic benchmark with pixel-level annotations, while the Potsdam-3 dataset offers a unique aerial, top-down perspective of urban scenes.
-The latter adds an additional challenge with its high-resolution images of buildings, roads, and natural elements captured over the city of Potsdam.
-For our qualitative analysis of LLM interpretation in the task of text summarization, as described in Section 4.3, we utilized the TL;DR dataset [49].
-The TL;DR dataset contains summary comparisons with human feedback collected by OpenAI.
-Each entry consists of a Reddit post, including its title, original content, and a human-generated TL;DR.
-Models.
-For all experiments in the vision tasks, we used CLIP’s image encoder ViT-B/32 [37].
-For interpreting text summarization, as described in Section 4.3, we used the Llama-2-7B language model [54].
-All experiments were run on 8 NVIDIA RTX 4090-24GB GPUs.
-4.1 Zero-shot Unsupervised Object Selection
-The upsampled relevance map $\tilde{S}_{i}^{(l)}$ can be converted into a binary segmentation mask using a threshold $\tau$, where the binary mask $M_{i}^{(l)}$ is defined as:
-$M_{i}^{(l)}[x,y]=\begin{cases}0,&\text{if }\tilde{S}_{i}^{(l)}[x,y]<\tau,\\ 1,&\text{otherwise}.\end{cases}$ (10)
-Here, $\tilde{S}_{i}^{(l)}[x,y]$ represents the relevance value at position $[x,y]$ in $\tilde{S}_{i}^{(l)}$, with $\tau$ as the threshold.
-When $\tilde{S}_{i}^{(l)}[x,y]$ is below $\tau$, $M_{i}^{(l)}[x,y]$ is set to 0; otherwise, it is set to 1, marking the object region.
-Our findings indicate that as tokens propagate through the network, they refine their object representation while retaining the semantic meaning of their associated image patches performing Object Selection.
-Figure 5 visually illustrates this process.
-For a given patch token $\mathbf{x}_{i}$, the object it most strongly represents is denoted as class $k_{i}$, indicating that $\mathbf{x}_{i}$ predominantly corresponds to a region of the object belonging to class $k_{i}$ in the image.
-The latent token $\mathbf{z}_{i}^{(l)}$ generates a relevance map that highlights areas with higher values associated with class $k_{i}$.
-After applying a threshold, this map becomes a binary segmentation mask expected to exhibit a high Intersection over Union (IoU) with the corresponding class $k_{i}$ region in the image.
-An illustrative example is shown in Figure 2.
-To quantify alignment, we compute the IoU by converting the relevance map $S_{i}^{(l)}$ into a binary mask $M_{i}^{(l)}$ and comparing it with the ground-truth mask.
-We propose the Initial Token IoU (ITIoU) metric, which measures how well the relevance maps of input tokens align with their respective class masks.
-The ITIoU is calculated as:
-$\textit{ITIoU}^{(l)}(X)=\frac{1}{C}\sum_{i=1}^{C}\frac{1}{|\mathcal{P}_{i}|}\sum_{\mathbf{x}_{j}\in\mathcal{P}_{i}}\text{IoU}(M_{j}^{(l)},G_{i}),$ (11)
-where $C$ denotes the number of classes, $\mathcal{P}_{i}$ represents the set of tokens associated with class $i$, $M_{j}^{(l)}$ is the binary segmentation mask for token $\mathbf{x}_{j}$ within class $i$, and $G_{i}$ is the ground-truth mask for class $i$ in image $X$.
-The inner sum averages the IoU for tokens in $\mathcal{P}_{i}$ for each class, and the outer sum then averages across all classes.
-Using a threshold of 0.2, our ITIoU metric achieves an average score of 37.84 % on the COCO validation dataset and 39.51 % on the VOC dataset.
-The ITIoU metric provides a comprehensive evaluation by incorporating a weighted average across tokens in each class, enhancing the assessment of token alignment with their respective ground-truth labels.
-Refer to caption Refer to caption
-(a) TL;DR: I’ve had very bad luck with girls my whole (b) TL;DR: I still like her but my rational
-life and I don’t know how to get my confidence up. side says ”no, she is a trash person”.
-Figure 4: Visualization of Token Contribution Scores ($\lambda_{i}^{(l)}$) highlighting the relevance of context tokens in interpreting the summary.
-Each token is colored proportionally to its $\lambda_{i}^{(l)}$ value.
-These visualizations demonstrate the model’s ability to identify key semantic elements in the context for generating relevant summaries.
-4.2 Zero-shot Unsupervised Semantic Segmentation
-We benchmarked the segmentation capability of our method against several approaches in the literature on unsupervised segmentation.
-Notably, unlike other methods, our approach requires no additional training.
-Instead, it relies solely on a pre-trained vision transformer, which may have been trained on tasks unrelated to unsupervised segmentation.
-To evaluate our method, we used the Unsupervised mean Intersection over Union (U.
-mIoU) and Unsupervised pixel Accuracy (U.
-Accuracy) metrics, following standard practices in semantic segmentation research.
-In all experiments across datasets, as shown in Tables 1, 2, and 3, we set the cutoff distance threshold $\zeta=0.4$.
-Table 1: Comparison of unsupervised segmentation methods on the 27-class COCO-Stuff validation set.
-”W” indicates methods requiring additional training, while ”W/O” denotes methods without additional training.
-Method Training U. Accuracy U. mIoU
-IIC [21] W 21.8 6.7
-DINO [5] W 30.5 9.6
-PiCIE [9] W 48.1 13.8
-ACSeg [27] W - 16.4
-STEGO [18] W 56.9 28.2
-DepthG [46] W 58.6 29.0
-U2Seg [33] W 63.9 30.2
-ULTra (Ours) W/O 67.2 32.9
-Table 2: Comparison of unsupervised segmentation methods on the PASCAL VOC validation set.
-Method Training U. mIoU
-IIC [21] W 9.8
-MaskContrast [56] W 35.0
-Leopart [66] W 41.7
-TransFGU [62] W 37.2
-MaskDistill [57] W 42.0
-ACSeg [27] W 47.1
-ULTra (Ours) W/O 51.9
-Table 3: Comparison of unsupervised segmentation methods on the Potsdam validation set.
-Method Training U. ACC
-IIC [21] W 65.1
-DINO [5] W 71.3
-STEGO [18] W 77.0
-DepthG [46] W 80.4
-ULTra (Ours) W/O 74.6
-Our results in Tables 1, 2, and 3 highlight the effectiveness of the proposed method for unsupervised semantic segmentation.
-Unlike baseline approaches, which require additional training specifically tailored to segmentation, our method leverages a pre-trained vision transformer without any further fine-tuning.
-Despite this lack of task-specific training, our approach consistently outperforms other methods, demonstrating adaptability across diverse datasets.
-4.3 Interpreting LLMs in Text Summarization
-In this section, we examine how our interpretability framework can be applied to text summarization tasks to uncover the underlying mechanisms and intent of LLMs.
-By visualizing the regions of the input context that an LLM prioritizes while interpreting a given TL;DR summary, we gain deeper insights into the model’s behavior and decision-making processes.
-As shown in Figure 4, these visualizations allow us to pinpoint the most influential regions of a textual input prompt in generating concise and relevant summaries.
-For the experiments, we used a Supervised Fine-Tuned (SFT) version of Llama-2-7B trained on the UltraFeedback Binarized (UFB) dataset [12].
-Additionally, we aligned the model to the text summarization task on the TL;DR dataset using the Direct Preference Optimization (DPO) method [38] for 1,000 iterations, with a learning rate of $5\times 10^{-6}$ and $\beta=0.5$.
-To validate our framework, we selected the preferred response (TL;DR) of each sample in the dataset, denoted it as $y$, and used it as the summary of the context $x$.
-In this task, we concatenate the context $x$ and the summary $y$ with a separator token.
-After feeding this input to the model, we compute the relevance scores of the TL;DR tokens with respect to the context tokens.
-We then average these scores for each token in $x$ to obtain a scalar value, referred to as the Token Contribution Score, $\lambda_{i}^{(l)}\in\mathbb{R}^{+}$, which highlights the contribution of each context token in interpreting the summary $y$ with respect to the context.
-This provides visual evidence of how the model processes the context text and identifies key semantic elements relevant to producing the summary $y$.
-Accordingly, $\lambda_{i}^{(l)}$ is computed as:
-$\lambda_{i}^{(l)}=\frac{1}{|y|}\sum_{j=1}^{|y|}S_{j+|x|}^{(l)}[i],\quad\forall i\in\{1,\cdots,|x|\},$ (12)
-where ||$\cdot$|| ⋅ | denotes the number of tokens in the text.
-(a) Original Image (b) Explanation Relevency Map
-Refer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-Figure 5: An example illustrating the model’s decision-making process across layers.
-Moving from left to right corresponds to deeper layers in the network.
-The first row corresponds to the CLS token, while the second, third, and fourth rows represent three different tokens, highlighted by red squares.
-(a) COCO-Stuff
-(b) PASCAL VOC
-Figure 6: Ablation study on two evaluation metrics across layers.
-These plots demonstrate a progressive improvement in semantic segmentation performance in the deeper layers of the transformer model.
-This enhancement is attributed to latent tokens capturing more meaningful segment structures, resulting in increasingly accurate and refined semantic representations.
-In example (a): semantically significant words such as ‘relationship’, ‘experience’, ‘rejection’, and ‘never’ are prominently highlighted, reflecting the model’s interpretation of the person’s struggles with relationships and feelings of rejection.
-Additionally, the highlighting of the question at the beginning of the context ‘How do I stop feeling bad…’ suggests the model recognizes the presence of uncertainty and a request for guidance, which is encapsulated in the summary as ‘I don’t know.’
-In example (b): $\lambda_{i}^{(l)}$ scores reveals the model’s focus on words such as ‘feelings’, ‘hate’, ‘disappoint’, ‘love’, and ‘like’, which correspond to the person’s mixed emotions toward their girlfriend, as described in the summary.
-The apparent contradiction between ‘love’ and ‘trashness’ in the summary seems to be derived from these highlighted terms, suggesting the model understands the conflicting emotions present in the text.
-Furthermore, the emphasis on ‘character’ aligns with the summary’s judgmental tone, suggesting that the model interprets this word as indicative of an assessment of personality traits and behaviors.
-This token-level analysis can serve as a valuable tool for effective supervision in future research, particularly for developing interpretable fine-tuning and alignment methods like Reinforcement Learning from Human Feedback (RLHF) [10, 49, 35] and Preference Optimization [38, 3, 14], where understanding model behavior and intent is essential.
-It may also provide useful insights in related areas, such as Chain-of-Thought [8, 24, 43] and Theory-of-Mind [60, 23, 61], which seek to make the reasoning processes and intentions of LLMs more transparent.
-Ultimately, our framework offers insights into fundamental questions in NLP, such as: ”How is the model thinking?” or ”What is the underlying intent behind the model’s generated response?”.
-4.4 Ablation Study on the Effect of Layer Depth in ViT Token Understanding
-In this section, we analyze the impact of depth on our model’s interpretability and segmentation performance, providing insights into the contribution of each layer.
-As anticipated and observed in Figure 6, deeper layers generally carry more semantic significance.
-However, the contribution diminishes in the final layers, suggesting that a depth of around 13 layers might be more than sufficient for the ViT to effectively comprehend image content.
-This finding implies that even fewer layers might achieve comparable results, potentially reducing computational costs without compromising performance.
-We observe an intriguing behavior in the initial layers, where performance initially declines before improving.
-This phenomenon is also visually evident in Figure 5, where the attention maps in the first layer appear to focus on the entire image.
-This suggests that, initially, the token examines the image as a whole before selectively gathering information from tokens with similar characteristics.
-5 Conclusion
-This paper introduces a novel interpretability framework that provides valuable insights into the decision-making processes of Transformer models.
-The framework is based on the input interpretation provided by each token in an arbitrary layer.
-By aggregating the interpretation of all tokens in the same layer, the framework enhances understanding of the Transformer’s behavior at the layer level.
-The richness of this understanding enables zero-shot unsupervised semantic segmentation, where our method achieves state-of-the-art performance with an accuracy of $67.2\%$ and an mIoU of $32.9\%$ on the COCO-Stuff dataset, and an mIoU of $51.9\%$ on the PASCAL VOC dataset for the unsupervised zero-shot version.
-Beyond vision tasks, the framework is not limited to vision models and can also be applied to interpret the behavior of large language models (LLMs) in tasks such as text summarization.
-This highlights the versatility and broad applicability of the proposed framework across various domains, offering potential for future work to simultaneously demystify the multimodal understanding of Transformers.